### v1

# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

Abstract.
We conducted co-design workshops with 23 participants (ages 15–24) to explore how youth envision an ideal remote social connection. Using the Fictional Inquiry (FI) method within a Harry Potter-inspired narrative, we found that youth perceive a disconnect between platforms labeled as “social media” (like Instagram) and those where they actually experience meaningful connections (like Minecraft or Discord). Participants envisioned an immersive 3D platform that would bridge this gap by prioritizing meaningful social connections, enabling time that feels well spent through presence and immersion, natural individual expression, intuitive social navigation that leverages physical-world norms, and playful, low-stakes opportunities for gradual friendship development. We introduce the design framework of spatial integrity, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. The FI method proved effective in generating innovative ideas while empowering youth by fostering a sense of hope and agency over the future of social media through their own design contributions.
social media; fictional inquiry; youth; design
conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3ccs: Human-centered computing Collaborative and social computing
1. Introduction
Most youth today use social media daily and spend significant time on platforms such as Instagram and TikTok (Anderson, 2023). However, these same youth often express deep dissatisfaction with how they spend this time—not because they want less social connection, but because they desire more meaningful online interactions (Kim et al., 2024c). They often feel constrained by the design patterns of archetypal social media platforms, which prioritize uninteresting and attention-grabbing content consumption over genuine connection (Landesman et al., 2024). This disconnect has contributed to broader societal concerns about social media’s impact on youth well-being (Oth, 2023).
Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide: 41 U.S. states suing Meta (Lima-Strong and Nix, 2023), proposals for “warning labels” on social media use (Murthy, 2024), restrictive legislation like the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act” (Oth, [n. d.]), and outright bans on social media use for those under 16 in countries like Australia (Kim, 2024). The EU’s GDPR (gdp, [n. d.]) similarly limits teen access.
However, such alarmist narratives oversimplify a complex reality: the effectiveness of these narratives and policies (as well as the causal role of social media in youth mental health) remains unclear. Despite recognizing the issues raised in these narratives, the youth report more positive than negative personal experiences (Vogels, 2023), even as they acknowledge potential risks to others (Vogels, 2023). Rather than perpetuating fear-based narratives or attempting to completely rebuild existing platforms, researchers and teens themselves emphasize the importance of learning to navigate social media effectively (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025). Furthermore, such alarmist responses and outright restrictions may be misguided—they impede the development of resilience (Wisniewski, 2018b) and induce extreme and unhelpful anxiety and mistrust without fostering constructive actions (Kim et al., 2024a), ultimately leading to lower psychological well-being associated with social media use (Lee and Hancock, 2024).
This misalignment suggests that the solution lies not in the wholesale rejection of social media or attempts to limit youth access, but in expanding our understanding of what social platforms can be. Rather than trying to “fix” existing platforms or promote fear-based narratives, we need to broaden our vision of what social media could be. To explore alternative visions for social technologies that prioritize meaningful connection, we employed the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) in co-design workshops with 23 participants (ages 15-24) experienced in 3D gaming environments (e.g., Minecraft (min, [n. d.]), Roblox (Roblox Corporation, [n. d.])). Participants were asked to envision a social media platform for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This magical context freed participants from current platform constraints while maintaining familiar social dynamics.
Our findings reveal two important patterns in how youth view and experience online social platforms. First, they make a clear distinction in their minds: platforms like Instagram are what they consider “social media”—where they spend significant time but rarely form meaningful connections. In contrast, they do not typically label platforms like Discord or Minecraft as “social media,” even though these are where they actually develop deep friendships and find their time meaningful. Second, because youth associate “social media” primarily with platforms like Instagram and its limitations, they become skeptical of social media’s overall potential to support authentic relationships. This narrow definition of social media, shaped by their experiences with Instagram-like platforms, leads them to dismiss the entire category as inherently flawed.
Participants identified several key elements missing from archetypal social media platforms: (1) emotional connection through presence and immersion, (2) diverse forms of individual expression that feel more natural, (3) intuitive social navigation—ranging from privacy to relationship development—that enables them to leverage norms from the physical world, and (4) playful, low-stakes opportunities to develop friendships. Although participants expressed deep dissatisfaction with the time they currently spend on archetypal social media platforms, they felt that spending the same amount of time on platforms designed to prioritize meaningful connections would feel valuable.
We contribute empirical findings on what youth perceive as the most significant shortcomings of current social media platforms, as well as what they desire to experience on social media. Additionally, we propose the concept of spatial integrity for social interactions in online spaces, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. These dimensions include spatial presence (the ability to understand the space and identify who is present), spatial composition (the meaning and significance of the space), spatial configuration (the organizational structure of the space), and spatial depth (the extent of information that can be perceived within the space). Furthermore, we demonstrate that the FI method, in the context of social media design, not only prompted creative idea generation and enabled participants to transcend current social media design patterns but also empowered them to envision the positive potential of social media, leaving them feeling more hopeful about its future possibilities.
2. Related Work
2.1. Defining Social Media and its Role in Young People’s Lives
Social media does not yet have a widely agreed-upon definition. There is a range of interpretations, from broad (e.g., “any interactive communication medium that enables two-way interaction and feedback” (Kent, 2010)) to specific (e.g., “a different kind of Internet-based applications which build the ideological and technological foundations of Web 2.0, and allow users to create content and exchange it with other people through the Internet” (Kaplan and Haenlein, 2010)). While some platforms, such as Instagram (All, [n. d.]) or Facebook (fac, [n. d.]), are indisputably considered social media, others, such as YouTube (you, [n. d.]) and Discord (discord page, [n. d.]), are less archetypal. There are also different types of social media, and even such classifications have multiple interpretations (Kaplan and Haenlein, 2010; Sharma and Verma, 2018; Scott, 2015). Some major categories include social networking sites (i.e., “web-based services that allow individuals to (1) construct a public or semi-public profile within a bounded system, (2) articulate a list of other users with whom they share a connection, and (3) view and traverse their list of connections and those made by others within the system” (Boyd and Ellison, 2007)) and blogs (i.e., “a way to share that love with the world and encourage an active community of readers who comment on the posts of the author” (Scott, 2015)).
The range of definitions and classifications of social media reflects the wide array of user needs that social media addresses. From a uses and gratifications perspective (Katz et al., 1973), social media supports a variety of utilities, ranging from social interactions to information seeking and entertainment (Whiting and Williams, 2013). Even within such broad categories of uses and gratifications, there are diverse ways these needs are pursued. For example, within social interactions, there may be a focus on bridging social capital and weaker ties versus bonding social capital and stronger ties (Phua et al., 2017). Further, even on a single platform—Instagram, for instance—users engage in a range of activities, from photo sharing for self-promotion and disclosure (Menon, 2022) to more passive consumption of media for entertainment. In other words, Instagram may serve as a personal branding tool for some while functioning as a way to pass the time for others.
Despite the prevalent narrative that social media is harmful, the experiences of youth with these platforms are more nuanced. They face challenges such as social comparison (Yau and Reich, 2019), significant privacy concerns (Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014), pressures to curate their online presence (Yau and Reich, 2019), and feelings of meaningless time spent (Lukoff et al., 2018), along with more subtle discomforts (Landesman et al., 2024). However, these platforms also offer meaningful benefits. For instance, marginalized populations often derive substantial support from these spaces (Li et al., 2023; Bates et al., 2020), and social media serves as an important medium for identity development and self-expression (Davis, 2012; Lee et al., 2022). Given this complex relationship between youth and social media, researchers advocate for strategies that focus on building resilience and equipping teens to navigate these experiences effectively rather than solely aiming to restrict or mitigate harm (Wisniewski et al., 2012b; Wisniewski, 2018a; Agha et al., 2023).
2.2. Social Technology Design Explorations in HCI
HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features. Social media platforms such as BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression. This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction. However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
At the feature level, efforts to support meaningful communication include designs that encourage effortful communication (Liu et al., 2021; Zhang et al., 2022) and those that foster feelings of genuine connections (Stepanova et al., 2022). While these features are important, social media experiences are often holistic, involving a complex interplay of multiple design aspects and user behaviors. As a result, such features alone cannot fundamentally alter the overarching narratives or perspectives surrounding social media.
Beyond traditional social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions. Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments. Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions. However, these platforms often are not generally recognized as social media in the traditional sense.
What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear—especially when considering the possibilities offered by emerging technologies such as spatial computing platforms. Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
2.3. Designing Beyond Doom Narratives
Social media and design fixation. Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design. While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking. This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
Social media design is no exception. Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps. However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)). These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
The Fictional Inquiry (FI) method. FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios. These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings. The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints. While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
According to Dindler and Iversen (Dindler and Iversen, 2007), implementing this method involves three key steps: first, clearly defining the inquiry’s purpose, whether it’s for staging design situations, exploring future ideas, or driving organizational change; second, developing an appropriate narrative that participants are comfortable with while being distinct enough from current practices to encourage new thinking; and third, creating a compelling plot that introduces tension or conflict within the narrative to motivate specific workshop activities. Unlike traditional role-playing approaches, FI encourages participants to maintain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities while drawing from their personal motivations and capabilities, all while circumventing typical societal constraints.
3. Method
We conducted co-design workshops with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms. All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n. d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers. During the workshop, we used the FI method to scaffold participants in generating design ideas.
3.1. Material: The Narrative Framework
Figure 1. Miro board mapping real-life relationships to Hogwarts-themed groups, designed to immerse participants in the Hogwarts-inspired FI narrative.
[…]
This is a visual chart designed to organize relationships and connections in the Harry Potter universe. There are three columns of labeled sections. In the middle, there are four sections, each representing one of the Hogwarts houses: Gryffindor (red), Hufflepuff (yellow), Slytherin (green), and Ravenclaw (blue). Each house box has its crest. To the left, there are two smaller boxes. The top-left box is labeled “Best Friends / Romantic Relationship(s),” featuring small images of Harry Potter, Hermione Granger, and Ron Weasley, with a space for a name or nickname. Below this, there is a smaller box labeled “House Elf,” featuring a picture of Dobby holding a glowing object and an area for a name. To the right, there are two more boxes. The top-right box is labeld “Family,” containing images of Sirius Black, James and Lily Potter, and the Dursley family. Below this, another box is labeld “Profs & Other Adults,” displaying images of Severus Snape, Dumbledore, Professor McGonagall, and Hagrid. At the bottom, there are three large sections. The leftmost section is labeled “Muggles” with a blank space. The middle section, labeld “Students from Other magical Schools,” displays crests from Beauxbatons, Durmstrang, Ilvermorny, and Castelobruxo. The rightmost section, labeled “Quidditch Team Members,” features an image of a Quidditch team in red robes. Each section contains an empty space for names or details to be added.
For our study, we selected the Harry Potter world as the narrative framework for our FI method. The Harry Potter series follows the journey of an orphaned boy who discovers he is a wizard and, as a result, is invited to attend “Hogwarts School of Witchcraft and Wizardry,” where he forms deep friendships and gradually uncovers his destiny as he clashes with the story’s antagonist.
Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI workshops. The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement. The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells. Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies. Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building. The characters’ experiences or life transitions—which parallel real-world moves between schools or locations—allow for exploring how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
3.2. Procedure
The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI workshop method. The interviews averaged around 90 minutes. We began the interview by asking which 3D games they had played and what they liked about those experiences. We also asked questions about Harry Potter, such as their favorite book in the series. This served to break the ice and help them start immersing themselves in the Harry Potter world.
Following this introduction, we presented the FI narrative: participants were asked to envision ways to support social connection within the Hogwarts world as a non-Wizard (i.e., “Muggle” (mug, [n. d.])) trying to stay connected with everyone across Hogwarts and the Muggle world. We then shared a Miro (mir, [n. d.]) board (Figure 1) where participants could sort their friends and family into categories inspired by the Harry Potter series, such as fictional houses (e.g., Gryffindor, Hufflepuff), teams (e.g., Quiddich Team members), and old friends (e.g., muggles), reflecting fictional houses, teams, and communities from the story. This activity encouraged them to think of their relationships within the Harry Potter context and was intended to further enhance immersion in the fictional design space. Next, we asked general questions about their experiences with social media, focusing on what they found meaningful or meaningless and their vision for an ideal social media platform.
During the design phase, participants were prompted to describe the capabilities their ideal magical social media would have and what aspects of their lives they would choose to share. Throughout this phase, we sought to keep participants immersed in the narrative, referring to existing platforms as “Muggle social media” and maintaining consistent engagement with the Hogwarts setting, as outlined by the fictional inquiry guidelines. Finally, we concluded by asking participants to reflect on how the social media they envisioned for Hogwarts compared to Muggle social media. The semi-structured interview protocol is available in Appendix A.
3.3. Participants and Recruitment
We used a combination of purposive and convenience sampling to recruit participants for our study. The first and second authors shared the recruitment materials (that included a link to a screener survey) on their social media accounts, while the first author also distributed flyers across our institution’s campus.
Participants were required to be between 15 and 24 years old, in accordance with UN (United Nations, [n. d.]) and WHO (who, [n. d.]) definitions for youth, and to consent to be video or audio recorded during a 90-minute interview. Video recording was requested following the FI method guidelines, enabling us to analyze gestures if necessary. We included both teenage and young adult participants (ages 15 to 24) to capture perspectives from distinct life stages, as these groups often navigate significant transitions and share a focus on identity exploration, self-presentation, and relationships.
Participants were also required to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox). This last criterion was based on the assumption that individuals familiar with 3D games would likely have experience with 2D platforms as well, equipping them with a broader set of cognitive tools for creative idea generation. By contrast, extensive experience with 2D platforms alone does not necessarily ensure the ability to conceptualize or engage with spatial technologies effectively. While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
Table 1. Demographics Summary. Abbreviations: Familiarity with Harry Potter - Expert (“I remember every detail about the plot, characters, and magical aspects (like how the castle moves, spells, magical creatures, etc.).”), Proficient (“I remember most details …”), Familiar (“I remember some details …”). Social Media Platforms – IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace). Reasons for Using Social Media – Connect (Connecting with friends and family), Network (Networking for professional purposes), News (Staying updated with news and current events), Share (Sharing content (photos, videos, etc.)), Community (Joining communities or groups with shared interests).
PID Familiarity with HP Social Media Platforms Reasons for Using Social Media Gender Age Ethnicity
P01 Proficient IG, DC, BR, RD Connection, Network, Entertainment, Share Woman 19 Asian or Asian-American
P02 Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI Connection, Network, Entertainment, Share Woman 20 Asian or Asian-American
P03 Proficient IG, DC, TT, X, RD, TH, NP Connection, News, Entertainment, Share, Community Woman 18 Black or African-American
P04 Expert IG, DC, SC, TT, BR, X, RD Connection, Entertainment Woman 16 Asian or Asian-American
P05 Proficient IG, DC, BR, RD, FB, TH Connection, News, Entertainment, Community Woman 18 White
P06 Familiar IG, DC, SC, TT, RD, FB Connection, News, Entertainment, Community Non-binary 18 White
P07 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, News, Entertainment, Share Woman 17 Hispanic or Latino, White
P08 Proficient IG, DC, SC, RD, FB Connection, Entertainment, Share, Community Woman 19 White
P09 Proficient IG, DC, TT Connection, Network, Community Woman 21 Asian or Asian-American, White
P10 Proficient IG, DC, SC, TT, BR, RD, TH Connection, News, Entertainment, Share, Community Non-binary 17 Hispanic or Latino, White
P11 Familiar IG, DC, TT, X, RD, FB, TH Connection, News, Entertainment, Share Woman, Non-binary 19 Native American or Other Pacific Islander, White
P12 Familiar IG, DC, SC, TT, X Connection, Network, News, Entertainment, Share Man 19 Asian or Asian-American
P13 Familiar IG, DC, SC, TT, BR, X, RD, FB, TH Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P14 Expert IG, DC, SC, TT, BR, X, FB Connection, News, Entertainment Woman 18 Hispanic or Latino
P15 Proficient IG, DC, SC, BR, X, RD Network, News, Share, Community Woman 20 Asian or Asian-American
P16 Familiar IG, DC, SC, BR, X, RD Connection Man 18 Asian or Asian-American
P17 Expert IG, DC, SC, TT, X, RD, TH Connection, News, Entertainment, Share Woman, Non-binary 17 Hispanic or Latino, White
P18 Familiar IG, DC, SC, TT, X, FB Connection, News, Entertainment Woman 23 Asian or Asian-American
P19 Expert IG, DC, SC, TT, X, RD, FB Connection, Network, News, Entertainment, Share, Community Self-described, GenderFluid/Trans Man 19 White
P20 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, Share, Community Woman 17 Black or African-American
P21 Expert IG, DC, SC, RD Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P22 Familiar IG, DC, SC, TT, X, RD Connection, News, Entertainment, Share, Community Woman 19 White
P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
A total of 84 individuals completed the screener survey. We reached out to the 48 respondents who met the recruitment criteria, and 23 participants provided consent and completed the interview procedure. The mean age of the participant group was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) identifying as women, 3 (13.0%) as men, 2 (8.7%) as non-binary, and 3 (13.0%) that were gender fluid. The full list of participant information is available in Table 1.
3.4. Data Analysis
We analyzed the interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides a systematic yet adaptable framework for interpreting qualitative data. The method’s independence from pre-existing theoretical frameworks made it particularly suitable for our study, given our study’s exploratory focus on youth-envisioned social media design through the FI method. It enabled us to effectively capture and analyze the diverse perspectives shared by our participants.
Our analysis began with the first two authors independently coding the same two transcripts line by line. During this phase, the codes were descriptive and closely adhered to the data. The authors convened to discuss and resolve any discrepancies in their coding. Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024). After four coding iterations, we reached saturation in our codes. Using this final set of codes, the first author proceeded to code all the transcripts.
3.5. Ethical Considerations
Given J.K. Rowling’s controversial statements regarding transgender individuals (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in our recruitment material and screener survey and reiterated our position during the interview procedure.
4. Results
Participants articulated frustrations with current social media and envisioned Social Media at Hogwarts (SMH) as a transformative 3D platform prioritizing authentic connections with close friends. They proposed customizable personal spaces, communal areas fostering meaningful interactions, and immersive shared experiences like gaming or walking together. They suggested features that allow users to share expressively, such as memory sharing and real-time updates via avatars or aesthetics alongside low-intensity interactions that offer presence and nuanced privacy. Participants prioritized user control over feed content, fostering closer relationships over unsolicited interactions. Table 2 summarizes this section, while Figure 2 visualizes the collective vision, and Appendix  3 outlines key features per participant.
Figure 2. An AI-generated illustration envisioning a platform whose design reflects a synthesis of participants’ ideas. It is a 3D, open-world social media world with personal spaces—where users mirror their identity and interests—and community areas that support shared activities and memories. These homes exist in a neighborhood-based spatial design, where proximity mirrors social closeness, enabling intentional navigation and serendipitous interactions rather than passive scrolling. The neighborhood also includes themed social rooms and community spaces for public gatherings and interest-based group interactions. Users can visit friends’ homes and explore shared memories and content, similar to a Pensieve, with selective access control over different areas, rooms, and objects, ensuring privacy and customizable levels of sharing. An intelligent pet guide assists in exploration, connection, and engagement, fostering personal expression and meaningful interactions beyond traditional social media.
[…]
A whimsical digital illustration depicting a fantasy-style virtual neighborhood at night, with warm glowing houses, cobblestone pathways, and floating orbs of light. The left side shows a village scene, where small, cozy houses are arranged in a grid, surrounded by gardens, fences, and street lamps. A blue glowing node and arrow highlight one particular house, linking it to the right-side image. The right side provides a closer view inside a warmly lit home, where a young girl in a sweater and dress interacts with a cake on a wooden table. A small holographic memory scene emerges from the cake with a floating, glowing pool of light, depicting a group of people joyfully gathered around the cake. A curious orange dog watches nearby. Shelves filled with books, a spiral staircase, and a glowing lantern enhance the cozy, magical atmosphere. The image suggests an interactive, immersive social media experience, where users can explore virtual homes and relive shared memories.
Table 2. Key Features for Envisioned Social Media from Participant Interviews
Desired Feature Key Idea Example Quote
Immersive Experiences Through Spatial Elements (Section 4.1) “In 2D, you’re just observing… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” (P01)
[Section 4.1.1]
Bringing Realism and Presence Through 3D
Emphasized need for physical presence and embodied interactions beyond flat screens “You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen.” (P06)
[Section 4.1.2]
Personal Spaces and Community-Centered Landscapes
Desired personal homes and neighborhoods that mirror real-world social structures “You each get your own house… anything you want.” (P03)
[Section 4.1.3]
Holographic Display and Virtual Reality
Sought more natural and immersive interfaces beyond traditional devices “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” (P01)
Organic and Intentional Social Interactions (Section 4.2) “The antonym to pressure, stress-free, community-oriented, wholesome, accessible, and accepting.” (P09)
[Section 4.2.1]
Flexible and Safe Spaces for Meaningful Conversations
Valued direct conversations in protected spaces “Actual conversations for real connections, individualized chatting… to get to know people… rather than just posting on a thread.” (P22)
[Section 4.2.2]
Community Gathering Places
Desired spaces for shared activities and casual encounters “a little card game area, shared experiences, something fun” (P06)
[Section 4.2.2] (1)
Activity Spaces for Shared Experiences
Valued collaborative activities and shared projects “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” (P06)
[Section 4.2.2] (2)
(Public) Third Places for Casual Encounters
Desired spaces for serendipitous interactions “You strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” (P03)
[Section 4.2.3]
Ambient Co-Presence
Valued low-intensity connections through shared virtual spaces “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.” (P19)
[Section 4.2.4]
Effortful and Intentional Navigation
Saw physical movement as a meaningful way to engage “Ideally, SMH would have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you.” (P05)
Expressive and Lower-Stakes Sharing (Section 4.3) “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” (P01)
[Section 4.3.1]
Expression Through Personal Area and Avatar Customization
Wanted customizable avatars and environments “decorat[ing] a house or customizing an avatar allows others to get to know more about my personality… through my style.” (P16)
[Section 4.3.2]
Real-Time Sharing: Capturing States and Changes
Desired ways to share real-time states “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that” (P04)
[Section 4.3.3]
Sharing Memories: Multisensory and Emotional Experiences
Wanted to share memories in an immersive and clearer way “jumping into a pool or taking the first bite of an ice cream sundae… from your brain, pull it out, and just be able to share that” (P04)
Granular, Intuitive, and Fun Privacy Mechanisms (Section 4.4)
“If people are posting more like their genuine real life and their experiences, privacy would be a little bit more complicated.”
[Section 4.4.1]
Space-Based Privacy
Wanted privacy boundaries based on spatial metaphors “A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.” (P13)
[Section 4.4.2]
Playful Privacy
Desired engaging privacy mechanisms with elements of fantasy “Secret knock or an object that reveals a secret passage to… secret rooms.” (P11)
[Section 4.4.3]
Contextual Privacy
Wanted context-aware privacy settings “Each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” (P14)
[Section 4.4.4]
Invisibility
Sought robust blocking and visibility control “I don’t wanna be able to tell that the person I blocked is even around me.” (P22)
[Section 4.4.5]
Age-Appropriate Spaces
Emphasized need for age separation “People who are 50 should not be engaging with people who are 13.” (P21)
Refocusing Interaction Priorities for Meaningful Engagement (Section 4.5) “We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page.” (P19)
[Section 4.5.1]
Intentional Content Consumption
Desired more control over content exposure “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor.” (P09)
[Section 4.5.2]
Reduced Celebrity and Commercial Influence
Wanted focus on real friends over celebrities “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” (P07)
[Section 4.5.3]
Spatial Representation of Relationships
Sought physical distance to represent emotional closeness “The top 7 people you’re closest to would be the 7 houses in your cul-de-sac area.” (P15)
[Section 4.5.4]
Intentional and Meaningful Use of Time
Valued active engagement over passive consumption “Anytime not scrolling and instead talking to people feels worthwhile, even if it’s just light conversation.” (P06)
4.1. Immersive Experiences Through Spatial Elements
Participants saw 3D and VR platforms as key to emotionally engaging, socially fulfilling interactions. P16 noted current social media (which we will refer to as “Muggle Social Media” or MSM for short) “does a good job at spreading information [but not] emotions,” arguing that “tight interpersonal interactions” require “talk[ing] to you and see[ing] your face,” which 3D and VR enhance. They envisioned SMH as a 3D open world with realistic, highly flexible social interactions, lifelike avatars, distinct public and personal spaces, and holographic interfaces that convey presence.
4.1.1. Bringing Realism and Presence Through 3D
Many found MSM “flat” or limited by “just a screen,” (P10) yearning for greater realism and freedom. P19 compared SMH to “social media like Minecraft” but broader in scope, supporting a “physical world through 3D graphics” (P19) with “more of a connection” (P06) and “an emphasis on communications.” (P06) P06 referred to the 3D massively multiplayer online game (MMO) Sky (sky, [n. d.]), where players can “talk to other people” or “sit down by a candle and chat” but wanted “more of a connection” than “just sitting there, talking to one random person.”
Navigating shared virtual spaces resonated with P03, who noted that actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” P06 similarly highlighted co-presence in 3D:
“it does feel more real… You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen… you’re sharing the chat room. But you’re not really sharing the space [like when] I’m playing Roblox. [In SMH] I’m sharing that space physically, even though it’s not really physical.” (P06)
4.1.2. Personal Spaces and Community-Centered Landscapes
Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world settings like homes, neighborhoods, and shared areas. P19 imagined “separate houses” with “a main house,” where users could “recreat[e] and decorat[e] your house.” P03 echoed this: “You each get your own house… anything you want” (P03). P21 expanded on the idea, seeing personal spaces as “subspaces” within a larger ecosystem.
Having a realistic landscape was essential for fostering connection and community for some. P06 described it as having “neighbors” or an “island next door” in a “town area.” This idea of community and proximity reflected participants’ desire to recreate the social structures of the real world in a digital setting. By mirroring the physical world, the landscape was also presumed to create a sense of place, grounding social interactions in a familiar and relatable context where users feel comfort and ownership. For example, P09 noted traversal in a corporal form would feel “fulfilling,” “calming,” and “reassuring.”
4.1.3. Holographic Display and Virtual Reality
A recurring theme was moving beyond traditional devices like phones and computers to embrace holographic and VR-based interfaces. Many participants felt mobile devices would seem cumbersome in a magical, immersive world. P01 noted, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting holographic screens, like “Jarvis in Iron Man,” or using spell-activated “panels” and “big screen[s].” P15 preferred VR glasses, calling them “more natural,” “easier,” and “more fun.” Other concepts included a “makeup mirror” as a portal, a “watch” that immerses users, or a “mirror glass” for video calls or summoning friends via the charm (i.e., accio).
The immersive nature of VR was seen as fostering deeper connections. P01 noted MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” They concluded 3D VR platforms are “more immersive and more awesome” (P03) than MSM, which is inherently limited “because it’s 2D.”
4.2. Organic and Intentional Social Interactions
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.”
In contrast, SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships.
4.2.1. Flexible and Safe Spaces for Meaningful Conversations
Many participants valued direct, synchronous conversations as the best way to build closer relationships. P22 wanted, “actual conversations” for “real connections,” preferring “individualized chatting…to get to know people…rather than just posting on a thread.” They envisioned “chat rooms you can walk into,” (P02) each dedicated to specific topics, such as a room for “high school stuff,” decorated with artifacts from shared experiences to spark conversations. These were seen as “protected space…where you can talk…without feeling unsafe or judged” contrasting the perceived lack of safety on MSM, where users “don’t know everything around it.”
P09 imagined a flexible, “Room of Requirement” (roo, [n. d.])-like system for creating customized rooms “on a moment’s notice,” reflecting the “purpose” of conversations. P07 also imagined that personalized chatroom aesthetics could make conversations feel more intimate, such as “photos of you together, inside jokes, or fun memories.” Interest-based spaces, like Discord (discord page, [n. d.]), were also valued for providing a “common point” (P13) for conversation and “a good talking point for introductions.” (P07)
4.2.2. Community Gathering Places
SMH’s envisioned community spaces blended intimate, private interactions with dynamic public areas, fostering shared experiences and casual encounters.
Activity Spaces for Shared Experiences. Participants saw shared activities as central to friendship-building. P06 envisioned “a little card game area” like Roblox, where “shared experiences” of “something fun” create “real memories” to be reminisced, “do you remember that time when…?” They added, “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” P12 valued collaborative “projects online,” such as “creat[ing] a house we love,” offering “a sense of self-achievement” and “teamwork” “instead of just like doom scrolling.” Sharing that experience together “on the same topic, on the same project” makes it “a better way to connect, instead of just… posting.”Even “mindless scrolling” became meaningful when shared. P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
(Public) Third Places for Casual Encounters. Although most of the interactions participants envisioned centered around friends, participants also emphasized public spaces for serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look! This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
4.2.3. Ambient Co-Presence
Participants valued the idea of low-intensity connections through ambient co-presence. P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background. They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations. P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.”
Music would facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They reflected on their experience:
“At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music. And it was one of the most awesome nights because we all knew in that moment that collectively we all enjoyed what we were hearing and it was just good. I don’t know how else to describe it. It was just nice.” (P19)
They added that music could also serve as a “conversation starter” while also providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
4.2.4. Effortful and Intentional Navigation
Physical travel was seen as a key way to make interactions feel more meaningful and intentional. While teleportation options like “spells” (P06) or the “Marauder’s Map” (P21) were seen as the convenient default, several wanted the option of being able to travel around the virtual world, as if they were taking a “nature walk” (P14). As P21 remarked, “Many times… the best memories happen during a walk.”
For some, the value of their redesigned interactions would come from the intentionality they would bring. P02 described engaging in SMH with deliberate purpose rather than habitual use, preferring a slower pace: “Teleporting would be nice, but not in this situation for me… I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” Similarly, P05 envisioned navigation in SMH as requiring focused effort:
“[Ideally, SMH would] have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you… you would have to walk through a town to get there. It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” (P05)
P14 highlighted how such intentional interactions on SMH would feel more “personal” compared to seeing “a random 2D Insta post that they posted 3 months ago.” Participants also noted that SMH’s intentional nature would amplify SMH’s focus on meaningful connections. P05 speculated, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
For others, the fulfillment came from the embodied experience of physical actions. P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.”
4.3. Expressive and Lower-Stakes Sharing
Participants emphasized that SMH should enable rich personal expression, allowing them to authentically and casually reflect their personalities and interests through memory-sharing and customization, from decorating dorm rooms to avatars.
4.3.1. Expression Through Personal Area and Avatar Customization
Participants wanted to create and personalize virtual environments and avatars that reflect their identities, preferences, and aspirations. Unlike MSM that only offers “two modes” (P10)— “light and dark mode[s]” (P10)—P21 envisioned a highly customizable world where users could create their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” They emphasized how this space could be shared with friends, “showing a little part of yourself.” P03 imagined a “dark academia” Hogwarts-inspired room filled with “so many bookshelves and… Victorian architecture,” to reflect their interests. P01 expanded on these ideas, envisioning a house with rooms “dedicated to each of your interests.” They added how that would enable them to showcase the different facets of their identities: “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.”
“3D avatar[s]” (P11) and profiles were another key way participants envisioned expressing themselves. P22 highlighted how platforms like Myspace (mys, [n. d.]) allowed for a level of customization that reflected individuality, describing how they enjoyed coding profiles to change “the font and really make it reflect me.” P01 also shared that there is an inherent “joy” in “just that process of creating it, and finding out who you are through that process.” Even just choosing a username was seen as a way to “convey to the world and your friends your own identity… which is a positive feeling usually.” Reflecting on Tumblr, P01 explained how such personalization is “mainly for you,” not just for others to see. P17 felt similarly, stating, “I don’t think it necessarily conveys personal aspects of yourself, but I think it will without you meaning it to.” As they put it, “your brain conjured that up, right? So I think that’s part of you.”
Participants highlighted how expression through choice feels authentic. For example, P01 felt that such “creative expression” could feel “a lot more personal” because “that’s not just like what life has given you… a reflection of your outlook on life.” P13 described this kind of indirect sharing as “mak[ing] social media an extension of everyday interaction,” allowing people to “gain the basic foremost information” about someone without requiring direct conversations. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
P10 explained that being in a 3D, VR environment amplifies this expressive nature, stating that 2D worlds feel more “compressed” and “able to see as much stuff as you can see in 3D.” P16 noted that in VR, others are “basically right in front of you when they’re talking,” making their “mannerisms… clearly easily observable.” They elaborated that one could learn a lot about someone based on “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.” This experience feels more authentic, as though meeting with “more real people… compared to the profile pictures.”
In addition to expressiveness, participants appreciated how these features support subtle and low-stakes sharing. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, valued the ability to “share things more subtly.” For instance, they explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style.” They found this form of self-expression to be “more justified” than posting, which felt “embarrassing” and “just doing this for other people.”
4.3.2. Real-Time Sharing: Capturing States and Changes
Another recurring theme was the ability to reflect real-time states through external artifacts in virtual spaces. P04 shared how they might opt into sharing updates on “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants envisioned that the blending of physical and virtual selves would allow for a more personalized medium of sharing. P01 imagined their avatar reflecting what outfit they’re wearing in the moment or what song they’re listening to, adding that this could help friends intuitively know “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 described how frequent updates to virtual rooms could reveal changes in personal interests, providing friends with “an insight into… what I’m thinking.”
Participants also discussed the ability to adapt avatars to reflect real-time emotions or actions. P21 described avatars that “read your emotions… showing [them] on your faces,” while P19 envisioned avatars mimicking real-life gestures. Music was frequently mentioned as a subtle and meaningful way to share current moods.
4.3.3. Sharing Memories: Multisensory and Emotional Experiences
A major theme was the ability to share memories directly with others, like the “Pensieve” (pen, [n. d.]). P04 imagined sharing experiences like “jumping into a pool or taking the first bite of an ice cream sundae” directly, creating more of an emotional connection. Sharing memories could also help reduce misunderstandings by enabling more accurate and nuanced expression, unlike MSM, where communication is mostly “just words.” (P03)
P07 analogically described memory-sharing as “screen sharing, but with your eyes.” P08 highlighted the practicality of memory-sharing, explaining how it could capture transient moments—like crossing a state line—without needing to pull out a camera. P03 highlighted how this would enable them to “just give” memories to their mom in real-time, instead of doing so “after the fact,” where they often found themselves missing the right opportunities.
P04 added that memory-sharing would foster more deliberate consumption of content:
“[SMH] feels more personal and interactive… if I could take more time in a less easy-to-consume format with what others choose to share.… On Instagram, when you click on someone’s page, you see all the posts at once. You don’t go one by one, taking time to read through every comment.… Here, you’d have to take more time to hone in on a memory…… It’s the satisfaction of consciously consuming content from someone you’re close to, being deliberate in viewing what they share, versus just quickly looking and being done.” (P04)
Some participants imagined that these memories would not just be sensory but should also include emotions and perspectives. P02 explained how sharing memories could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02) For some, memories weren’t just visual; P01 expressed a desire to “transmit smells through social media,” adding another sensory dimension to shared experiences. Given ethical concerns, many felt that only minimal editing, such as “cropping certain parts you don’t want to share” (P01) or versions “sped-up or slowed-down,” (P07) should be allowed.
4.4. Granular, Intuitive, and Fun Privacy Mechanisms
Participants felt that current social media platforms offer limited privacy options, often lacking flexibility and nuance. P16 noted, “It’s kind of a pain to organize, and… right now in Instagram you can put close friends… that’s about it.” They expressed a need for more granular controls, particularly in a platform like SMH, where users might share more authentic aspects of their lives. As P11 explained, “If people are posting more like their genuine real life and their experiences, [privacy] would be a little bit more complicated,” emphasizing the importance of having “flexibility” to control access.
4.4.1. Space-Based Privacy
Participants envisioned using spatial or visual cues to establish privacy boundaries, drawing on intuitions from physical spaces. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
“I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home? If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
P13 elaborated on how privacy norms in physical world contexts transfer to the SMH world, stating:
“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate. It’s where you spend most of your time, and it shows your personality the most. It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes—or 30 seconds—the way you can on Instagram. If you wanted to add someone there, all you’d have to do is remove a couple of things. It’s easier than physically changing a space.” (P13)
While managing visibility across multiple rooms might initially seem more complex, participants believed this approach would feel intuitive and natural. P11 suggested that spatial contexts could simplify privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.4.2. Playful Privacy
Participants envisioned engaging privacy mechanisms imbued with a sense of playfulness and fantasy. P11 proposed a system where houses or specific objects within personal rooms could incorporate elements like a “secret knock” or an object that reveals a “secret passage to… secret rooms” when interacted with in a specific way known only to an exclusive list of friends. Others suggested knowledge-gated access, such as a “secret handshake” (P21) with their friends. P02 drew inspiration from the Marauder’s Map, envisioning a system where users could only see accessible spaces unlocked by summoning secret “spells.” P05 likened playful privacy mechanisms to “a little Easter egg”, stating “There’s that little kid part of everyone” that makes the features enjoyable. They noted that these elements would feel “fun” and “novel” as opposed to MSM’s privacy settings, which often come across as uninspired or confusing, “like a wall of settings where sometimes the title of what the setting is doesn’t even make sense.”
4.4.3. Contextual Privacy
Participants reflected on the challenges of navigating context collapse on MSM (Marwick and Boyd, 2011), and hoped SMH would make contextual disclosure easier:
“On Muggle social media, you’ve got to create a whole new account. But I’m thinking for this app, it just telepathically knows what’s in Muggle area and what’s in Hogwarts area. So you don’t really have to go through the effort of keeping two lives. You can just walk through a portal. And you’re like a new person.” (P02)
P03 proposed a framework where every user has a “base” identity but could customize additional aspects of their representation. P22 imagined a contextual profile where “any group you’re in…you could change your avatar for that,” drawing inspirations Discord where “you can make different profiles for different servers.” P14 highlighted the importance of individual control, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.”
Participants also emphasized that contextual privacy goes beyond person-based access control. For example, they suggested varying access based on temporal factors. P14 proposed that “people who don’t follow you… can’t see past versions… I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” The idea of gradual openness was another suggestion:
“if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.” (P21)
4.4.4. Invisibility
Participants were critical of existing platforms for their lack of a robust blocking system. P22 noted that on Instagram, even with contact syncing off, “they might still suggest accounts based on the follow relationships,” exposing a Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they would rather avoid. They also emphasized the need to block someone thoroughly, including their “alternate accounts,” that “I don’t wanna be able to tell that the person I blocked is even around me.” P14 added that on Instagram, even private accounts allow “a random person” to see personal content such as Story highlights (Bojkov, 2019).
Many users expressed a desire for greater control over their visibility on social media, emphasizing features that would allow them to be effectively “invisible” unless they chose otherwise. P03 envisioned users having “an invisibility cloak” for situations where “you don’t want anyone to see you.” Similarly, P07 proposed a “super private” profile option where others “can’t even search… like you’re not even on the app.”
The ability to prevent unsolicited contact was another recurring theme. P22 suggested a model inspired by Discord, where “unprompted contact” happens only with people “you’ve already interacted with,” and others “can’t even be discovered unless you have something in common.” This approach resonated with P13, who felt strongly that “people should only be able to add you if you meet in person and exchange a QR code.” They reflected on an experience where they were able to find personal information about a waitress online based just on their name, noting “it’s too easy to find people on social media.”
4.4.5. Age-Appropriate Spaces
Participants emphasized the importance of creating age-appropriate spaces for minors. P21, for example, argued “people who are 50” should not be engaging with “people who are 13.” Age verification was widely seen as a critical first step. P16 suggested incorporating “biometrics” into the verification process, while P19 argued that users should “legally prove” their age. Once reliable age verification is in place, participants advocated for creating separate spaces for minors to which older users do not have access. Conversely, P22 proposed that minors should have restrictions, such as being unable to join general spaces like a “town square.”
4.5. Refocusing Interaction Priorities for Meaningful Engagement
Participants expressed a strong desire to refocus their social media experiences on meaningful connections and intentional interactions. This section explores four key themes that emerged from their responses: the need to prioritize genuine connections over algorithmic content recommendations, reducing the influence of celebrities and commercial content, using spatial metaphors to represent relationship strengths, and fostering more intentional and meaningful use of time. These themes reflect participants’ vision for social media platforms that better align with their desires for authentic social connection and purposeful engagement.
4.5.1. Intentional Content Consumption
Participants emphasized the importance of refocusing social media on genuine, personal connections rather than extraneous content. P19 criticized how platforms like Instagram have shifted away from its initial focus:
“We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page. We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
This sentiment was echoed by P22, who lamented that on Instagram, “half of the timeline is accounts I don’t even follow that were just randomly recommended to me.” They described this experience as “less keeping up with friends and more just getting random stuff shoved in my face.”
On a related note, a major source of frustration was the infinite scrolling design of current platforms. P11 explained, “There’s always more content on social media… this fuels the mindset that… I need to make sure I’m not missing anything.” P14 also criticized the endless content stream on TikTok, which exposes users to “clickbaity” and often harmful content that “doesn’t stimulate [users] at all” and “messes with [users’] attention span.” They expressed a desire to “be more proactive” and “genuinely use [their] brain” instead of “just scrolling mindlessly for an hour.” P09 envisioned a system like an elevator: “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor… Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” This sentiment was echoed by several participants who envisioned brain-controlled interfaces that would “automatically show [users] things [they] would really like” (P07) based on their immediate needs or interests.
4.5.2. Reduced Celebrity and Commercial Influence
Several participants advocated minimizing celebrity influence. P02 stated, “I don’t want celebrity stuff… I don’t want any celebrity drama,” while P01 suggested removing celebrities altogether if the platform prioritized personal enjoyment. If included, they proposed isolating them to a “celebrity zone.” Prioritizing posts from close friends was a key preference. P07 explained, “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH offering greater control over interactions and content, noting it would be “more designed for people you actually know… like your friends, classmates, or Quidditch team,” rather than celebrities or acquaintances. P19, similarly, envisioned that on SMH, they could “filter out for friends and family, and even then condense your circle a little bit more.”
Other participants expressed a need for control over how non-personal content is organized and placed on the platform. P04 suggested being able to “figure out how I want to arrange strangers’ spaces in my… platform.” They also wanted the ability to “prioritize things and feel more in control of how much information is fed to them,” imagining a system where apps and notifications are ranked by personal relevance. P18 saw ads as inevitable but wished they could appear in a separate “billboard” that users could skip.
4.5.3. Spatial Representation of Relationships
Youth envisioned SMH as a platform that reflects the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties. Participants appreciated this design as it prioritized interactions with closer friends, creating a sense of community and authenticity. P02 highlighted, “The fact that only my close friends have access to it… is the defining factor,” enabling users to focus on “better conversations with other people [that] already know who I am,” without the pressure to “advertise who I am” as they might feel on MSM.
Some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction. P03 imagined, “The people around you… would be people from the same house as you,” cultivating a shared sense of “togetherness” and “community building” through common spaces and interactions. Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them.
This design approach—aligning physical landscapes with emotional closeness—was seen as a way to streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.” Unlike MSM feeds, which participants felt prioritized engagement over relationship strength, SMH’s spatial arrangement allowed users to engage naturally with those closest to them. P09 further elaborated on how this physical representation could enhance navigation: “have your friends’ rooms, or like your kind of close friends list, and maybe you’re kind of living in a pod with them, so you can move between the rooms a lot quicker.”
4.5.4. Intentional and Meaningful Use of Time
Participants expressed a strong preference for digital experiences that feel meaningful and intentional, contrasting these with the negative feelings often associated with aimless social media use. P10 described how social media can feel “meaningless” when they “didn’t learn something,” while P18 reflected on the regret they feel after “mindlessly scrolling.” Similarly, P14 shared that scrolling through Instagram Reels while “literally just laying in [their] bed” feels like “wasting [their] time.”
In contrast, participants found interactions involving direct communication and engagement more fulfilling. P16 shared that “anytime not scrolling” and instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this sentiment, noting that social media use feels more meaningful when they are “actually talking to someone.” They explained that such interactions create a sense of “connectivity,” particularly when keeping in touch with people “not in close proximity.”
The distinction between passive and active engagement was another recurring theme. P10 explained that while doomscrolling does not make them feel “fulfilled,” playing games like Minecraft is different because it involves “progress” they can “save and come back to.” Similarly, P14 shared that the lack of connectivity on social media drives their “doomscrolling,” but they believe platforms that encourage interaction could reduce screen time by making time spent online more meaningful. P06 also note that they noted that they “don’t feel bad… most of the time [they] use Roblox,” as it involves engaging activities such as “playing games with friends” and “customizing spaces.”
Given such differences in how time spent is perceived, participants expected time spent on SMH to feel more meaningful than that on MSM. P04 noted that while MSM allows for quicker interpersonal interactions, it often leads to “doomscrolling.” In contrast, SMH would, for example, encourage “thoughtfully engaging with someone’s memories,” leaving them feeling “much more fulfilled.” P05, similarly, explained that they would not feel bad about spending more time on a platform like SMH because it involves “more effort and more time… a more intentional way of interacting.” P07 affirmed this, stating they would be willing to spend more time on a platform like SMH given interactions where they would find interactions “more enjoyable.”
4.6. Reimagining Social Media’s Potential
Given the nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what social media could be. Therefore, while they acknowledged potential challenges—such as the platform being overly immersive, addictive, or demanding—they largely viewed these downsides as worthwhile trade-offs for its many benefits. Overall, participants saw SMH as an inspiring alternative to existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional. They also appreciated the FI workshops for empowering them to reimagine social media and to be hopeful about its possibilities in ways they had previously thought impossible.
4.6.1. Fostering Hope for the Positive Potentials of Social Media
Participants saw SMH as a platform that offers better user experiences, departs from the repetitive structures of existing platforms, and inspires hope for what social media can become. For example, P06 shared how SMH combines several desirable aspects of social media in a cohesive way, creating an environment focused on meaningful interactions: “You get friends who you can talk to, keep up with, and stay connected to. And you get things you can customize. You kind of get the whole package there. I kind of like all of those here.”
Participants also appreciated how SMH offered a break from the monotony of existing platforms. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram. P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “fun” and “novel,” like a “breath of fresh air.”
Beyond features, participants expressed how SMH inspired hope for the future of social media. P10 pointed out that overcoming the design fixation around social media is crucial:
“The way that we can make social media better is if we expand the realm of what the definition of social media is. Instead of it being where someone just shares stuff, it can also be places where people connect, and people do this and that.” (P10)
Several participants noted that the FI workshop played a pivotal role in helping them gain a broader vision of social technologies and envision new possibilities for social media. P15 reflected on how their perspective evolved throughout the process:
“When we started, I literally was giving all these suggestions, which was completely the same thing which existed… I was repeating their concepts. So I think throughout this entire interview, I evolved my idea… actually thinking about what is new… So I think this has been a super hit. I like the entire conversation.” (P15)
They appreciated that SMH was not merely a “clone of the same thing… without actually adding anything else” but rather a concept that was “completely different… not even the same.” Similarly, P14 shared how the workshop shifted their initial skepticism:
“I don’t know what else you could do, because it’s online… so I really don’t know what else there could be to make it more meaningful.” (P14)
After engaging with the workshop, they felt empowered to imagine broader possibilities:
“I’m not really a tech person. So I don’t really think about [how to make social media better] that much. But I feel like this fake scenario… [helped] remind myself, ‘Girl, you have full control, this is no restrictions.’ After I got into that I… have a way bigger vision of what social media can become. (P14)
4.6.2. Celebrating Youthful Creativity and Play
SMH allows a playful and engaging social media experience, allowing users to immerse themselves in actively creative activities and self-expression. P06 appreciated this ability to personalize their experience on SMH, stating, “I think… everyone likes sharing parts of themselves.” P21 likened SMH to a “game” due to its creative freedom. They elaborated, “The creative freedom you have over the platform makes it feel more fictional than it is real.”
Participants found the overall experience of SMH to be youthful, fun, and deeply immersive, offering a space that feels both magical and meaningful. For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
“Yeah, [I would prefer the 3D version] actually 100%. I feel like,… the only reason I said 2D is because I felt like the need to be like a mature adult. But who cares? I feel like something that’s more enjoyable would be [the SMH].” (P14)
P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms. P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness. P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
4.6.3. Enabling Otherwise Inaccessible Experiences
SMH creates opportunities for users to engage in experiences and interactions that may otherwise be out of reach. P09 highlighted how individuals “experiencing stress” or those who may not have access to “really, really beautiful landscape and environments” could walk around and “admir[e]” the “calming” landscape of SMH. P16 envisioned SMH as a space where users could engage in activities like climbing a mountain in Nepal, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
P02 saw SMH as a platform that would “seem more like real life, but better.” They imagined being able to engage with friends in idyllic settings, such as “on the beach all the time.” Furthermore, they believed SMH could strengthen friendships through magical features. In their words, “Since people can’t read others’ emotions in real life, SMH would… be a way for you to connect with your friends on a deeper level.” P18 also appreciated how SMH’s 3D interface facilitated them to interact with “family across the world” in an immersive way. SMH would offer a sense of presence that goes beyond conventional video calls, enabling users to engage with loved ones “not just through a screen.”
5. Discussion
Figure 3. (a) Current youth perceptions of social media platforms and where they experience meaningful connections, highlighting the disconnect between platforms commonly recognized as “social media” versus where youth actually experience meaningful social connection. The SMH—youth-envisioned ideal social media—is represented by the overlap, combining both aspects. (b) A more nuanced viewpoint clarifies “social media” as an umbrella term for a range of platforms with different primary purposes—friendship-building, entertainment, status, information-seeking, tie maintenance, and more—rather than a few archetypal platforms.
[…]
This diagram is divided into two parts, offering insights into how youth perceive social media and how they categorize its purposes. On the left, there are two overlapping circles. The circle on the left represents platforms that youth identify as “social media,” such as Instagram, Facebook, and TikTok. The circle on the right illustrates platforms where youth feel meaningfully connected with friends, including FaceTime, Discord, Minecraft, and Roblox. In the overlapping area falls Social Media at Hogwarts (SMH) youth-envisioned ideal social media that combines the qualities of both categories. On the right, a Venn diagram shows potential categories of social media platforms by their primary purposes, such as “Entertainment-Centered,” “Status-Centered,” and “Information-Seeking-Centered.” “Social Media at Hogwarts” is grouped under “Strong Tie-Centered.”
5.1. “Social Media is 2D”: Youth Perspectives on What Constitutes Social Media
Public discourse increasingly portrays social media as fundamentally detrimental to mental health, friendship, and authentic social interaction. This narrative often equates social media with specific platforms like Instagram, TikTok, or X and the ways in which they happen to be designed today (Kim et al., 2024b).
Our study participants expressed similar perspectives, frequently describing social media as centered around superficial interactions that fail to foster meaningful connections. As illustrated in Figure 3(a), there is a disconnect between platforms that youth commonly identify as “social media” (such as Instagram and TikTok) and the platforms where they actually experience meaningful connections with friends (such as Discord or Minecraft). When envisioning their ideal social platform, participants described something that would bridge this divide—a space combining the deliberate pursuit of social connection they associate with traditional social media with the organic, engaged interactions they naturally experience in gaming and chat platforms.
This misalignment between expectations and experiences around social media is problematic because negative perceptions of social media use can diminish psychological well-being (Lee and Hancock, 2024; Kim et al., 2024a). Youth experience disappointment in seeking deep social connections on platforms where such interactions are scarce. As shown in Figure 3(b), social media serves diverse uses and gratifications—from information seeking and entertainment to status enhancement and network maintenance. Rather than maintaining an entirely negative view of social media confined to certain platforms, it is crucial to understand each platform’s distinct values and limitations.
For instance, critiquing Instagram for not facilitating authentic, deep friendships misses the point; while such connections might occasionally form there, it is not the platform’s central form of engagement. However, it is important to note that platforms aren’t static—their designs evolve, and user engagement patterns shift over time. Instagram, primarily used for personal branding, sometimes facilitates closer communication through private ‘finsta’ accounts. This fluidity in platform use underscores the importance of teaching (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025) youth to be more discerning about the connections they seek and what’s possible on each platform rather than enforcing anxiety-inducing narratives about social media’s harmful effects.
5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings confirm prior studies that while the youth are largely dissatisfied with social media and how they spend their time on archetypal social media platforms, they do not necessarily want to reduce their social media usage—rather, they seek platforms that facilitate deeper, more meaningful connections. Our participants identified four key elements that could transform social media from a space of passive consumption to one of genuine connection:
  1. (1)
    Presence and Immersion. Youth envision platforms that foster genuine emotional connection through real-time presence. While current platforms feel “flat” and emotionally disconnected, participants valued synchronous interactions that create a sense of co-existence. Even simple activities like sharing quick updates about eating ice cream or walking outside, or simply existing in a shared virtual space were seen as meaningful when done with immersive co-presence.
  2. (2)
    Supporting Individual Preferences. Youth desire more natural and diverse ways to express themselves. Participants envisioned platforms where they could show their personality through both active sharing and subtle signals, similar to how personality naturally emerges in physical spaces. From memory-sharing to customizing their virtual spaces and avatars, they want to express themselves spontaneously and casually rather than feeling constrained to carefully curated posts.
  3. (3)
    Intuitive Social Navigation. Meaningful social media experiences require design that mirrors how relationships naturally unfold. Youth value platforms that enable organic relationship growth through proximity and repeated low-intensity interactions, similar to how sitting in the same classroom gradually builds connections. They also seek designs that replicate the subtle, embodied norms of physical spaces—where spatial arrangements naturally signal degrees of privacy and social contexts, much like how a bedroom implies different privacy expectations than a living room.
  4. (4)
    Playful Engagement. Youth emphasize that meaningful connection often grows through joy and shared experiences, yet they find current platforms overly serious and high-stakes. They envision platforms that encourage whimsical exploration and casual interaction—where they can discover hidden objects, explore virtual landscapes, or simply hang out in a “third place.” This playfulness creates low-stakes opportunities for friendship development beyond the performative nature of mainstream social media.
These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams. Instead, youth express a clear desire for platforms that prioritize meaningful connection over superficial engagement. While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
Our participants envision environments where relationships deepen naturally through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement—challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
5.3. Spatial Integrity for Meaningful Social Connections in Online Environments
In the design of virtual environments that foster meaningful social connections, spatial cognition—the way we perceive and interpret space—remains critically relevant. Even in virtual spaces, human interactions are shaped by spatial affordances, mirroring many aspects of physical-world interactions. While virtual spaces need not replicate physical environments exactly—indeed, they can transcend physical limitations in some ways while falling short in others—certain spatial principles remain fundamental to how humans build and maintain relationships. This relevance is reflected in our pervasive use of spatial metaphors when describing virtual interactions, such as “chat rooms,” “cyberspace,” or “folders.” While some platforms like Discord and social VR applications successfully implement spatial design elements, mainstream social media platforms often lack these affordances. Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces.
  1. (1)
    Spatial Presence: Understanding the Space and Who’s There. Spatial presence refers to the ability to perceive and navigate within a virtual space, encompassing elements such as proximity/propinquity (knowing who is nearby), navigation (moving within the space), and awareness (understanding one’s position relative to others). While gaming platforms and social VR environments often excel at creating this sense of presence, mainstream platforms like Instagram lack basic spatial affordances, limiting users’ awareness and intentional navigation within the platform.
  2. (2)
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offers only transient content streams. Our participants emphasized the importance of dedicated areas—like communal lounges or study spaces—where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
  3. (3)
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity—a user’s feed” is meant to be personal space yet lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
  4. (4)
    Spatial Depth: How Much Information Can Be Perceived. Spatial depth determines how much information users can perceive in a space. 3D spaces enable higher perceptual density, allowing users to process information more intuitively (Kumar and Benbasat, 2004; McIntire et al., 2014; Tavanti and Lind, 2001). Current platforms, limited by 2D interfaces, restrict users to explicit actions like posting and commenting rather than supporting more nuanced forms of expression. Incorporating spatial depth could enable richer self-expression and more natural interactions.
These four dimensions of spatial integrity—presence, composition, configuration, and depth—are critical for designing virtual environments that support meaningful social connections. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram that we explored in Section 5.2. As spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).)
The concept of spatial integrity offers a framework for understanding why many social media platforms feel disconnected and unsatisfying. By thoughtfully incorporating these principles, we can design online environments that better align with human social-spatial cognition, addressing the limitations our participants identified and better supporting meaningful social interactions and the organic development of relationships.
5.4. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
Our interviews reveal how mainstream social media’s design fixations (Jansson and Smith, 1991) have shaped youth’s perceptions of what social platforms can be. Youth have come to equate social media with certain limiting characteristics: endless content streams, superficial interactions, and 2D interfaces focused on broadcasting posts. Interestingly, while they report meaningful social connections and positive experiences in other online social spaces like Minecraft or Discord, they view mainstream social media as inherently tied to emotionally disconnected experiences and superficial interactions. This cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The Hogwarts-inspired context provided fresh metaphors for exploring social spaces, privacy, and connection. Unlike traditional design workshops where youth typically propose incremental improvements to existing platforms, these sessions enabled participants to imagine entirely new possibilities for meaningful connection. More importantly, the workshops helped participants expand their understanding of what social platforms could be, empowering them to articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This study demonstrates how design methods can offer a generative path toward hope in domains where narratives of doom are prevalent.
5.5. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, potentially biasing them toward favorable attitudes toward incorporating spatial elements or online environments that center play. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research could empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him—including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.
6. Conclusion
Through co-design workshops employing the Fictional Inquiry (FI) method and leveraging a Hogwarts-inspired narrative, we explored the youth’s vision for an ideal social platform. The interviews revealed a disconnect between mainstream social media and the kinds of meaningful online connections young people seek. While platforms like Instagram have become synonymous with “social media,” youth find their most authentic social interactions happening elsewhere—in gaming environments, chat platforms, and other spaces that they don’t typically seek as destinations for social connection.
The participants collectively emphasized four key elements missing from current social platforms: emotional connection through real-time presence and immersion, natural forms of self-expression beyond carefully curated posts, intuitive social navigation that mirrors physical-world relationship development, and playful low-stakes opportunities for friendship growth. These findings challenge the widespread assumption that engaging youth requires addictive features or endless content streams. Rather, young people express a clear desire for time that feels meaningfully spent, imagining that social media platforms enabling meaningful connection through shared experiences and deliberate engagement would provide exactly that.
Importantly, this research demonstrates how moving beyond standard design patterns through methods like Fictional Inquiry can help youth articulate possibilities they might have previously dismissed as impossible within conventional social media paradigms. Rather than accepting current platforms’ limitations as inevitable, participants were able to envision radically different approaches to online social connection and feel empowered and hopeful about the positive potential of social media that they initially felt pessimistic about.
Further, we introduce spatial integrity as a framework for facilitating meaningful social connections in online environments, helping to understand why many social platforms feel disconnected and unsatisfying while pointing toward design principles that could better support meaningful online relationships. From 2D platforms to spatial technologies, our study highlights an opportunity and hopeful future for creating social experiences that work with, rather than against, our natural needs and intuitions for social connection.
Appendix A FI Workshop Protocol
Introduction
Welcome, and thank you for participating in our design workshop. Before we begin, I want to ensure that you’ve read the consent form I sent earlier. Have you had a chance to review it?
[Response]
Great! Just to reiterate, this session will be recorded for research purposes. Do I have your consent to proceed with the recording?
[Response]
Thank you. If at any point you feel uncomfortable or wish to stop the interview, please let me know.
General
Let’s start with some general questions about your experiences with 3D gaming and Harry Potter.
  • Which 3D games do you play and how long have you been playing them?
  • (Which is your favorite?) Can you describe the 3D elements or the spatial elements in that game a little bit, please?
  • Have you ever played in VR mode if that is available?
  • Any AR games you have played? Like Pokémon Go?
  • What Harry Potter books or movies have you watched?
  • Which is your favorite?
  • Which Hogwarts house would you put yourself in?
Narrative/Plot
As mentioned in the consent form, our goal today is to explore innovative ways to improve social media. We particularly invited 3D game players with a love for Harry Potter because we believe this will help us think creatively and “outside the box” about what an ideal, magical social media could look like, beyond the confines of a small 2D smartphone screen.
Don’t worry about being too creative—we’re here to guide you through the design process, not to evaluate you. We want to explore different design directions with your help, and we’ll be actively involved while also giving you the space to share your ideas.
To give our brainstorming some structure, we’ll start with a specific scenario to solve during this session. I’m going to ask you some guiding questions to help you address the scenario:
Imagine you are a student at Hogwarts. You have friends who are Muggle-born, friends from magical families, and even friends at other wizarding schools around the world. You’re also getting to know me, as I’m a new student you’ve just met at Hogwarts. Keeping in touch with everyone is tough because traditional Muggle social media like Instagram doesn’t really portray your wizard self very well. You don’t have a smartphone in the first place because Hogwarts people don’t need electronic devices, so you can’t really charge your smartphone.
One day, a brilliant Muggle-born student had an idea. They decided to use magic to create a magical social media where everyone—Muggles, Hogwarts students, professors, and friends from other magical schools—could all communicate and share their lives in the most ideal way. Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?
Character Map
Before we dive into the designing part, let’s first map out your relationships in your new life as a Hogwarts student. Use the Miro board link (Figure 1) to write down the names or nicknames of people that you would like to put in each section. For example, you might want to put “younger sister” or “Jane” under Family, some of your coworkers under Quidditch Team members, yourself under the house that you think you belong to, your older middle school friends as Muggles, etc.
Please note the two boundaries in the four rectangles for each house. You can place your closer friends in the inner box and not-so-close ones in the outer box. And you can put the same name in multiple boxes (or parallelograms) as you wish.
  • How do you currently communicate with or stay connected with each of these people via Muggle social media?
  • What are your biggest pet peeves/issues with Muggle social media that you’d like to solve with magical powers?
  • What moments do you feel bad/guilty about using social media, if at all?
  • When does time spent on social media feel meaningless/meaningful/fulfilling, if at all?
  • Pet peeves around privacy settings?
  • When do you feel most connected to your friends when interacting with them on social media?
  • What would a better social media be if you were to use adjectives to describe it? What does “better” mean to you?
Design
Now, let’s think about what this magical world social media might look like. Remember, there are no right or wrong answers here. We’re looking for your imagination and creativity, so feel free to think outside the box and have fun with it!
  • What key capabilities would this thing have for communicating with [GROUP]*?
  • If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)
  • What would you want to share about yourself on this ideal magical “thing” with [GROUP]?
  • If you had [GROUP’] join this “thing,” how might its form adapt to include this new group?
*: Iterate for 1) Romantic Relationships and Closest Friends, 2) Closer Hogwarts Friends, 3) Muggle Friends, 4) Friends at Other Magical Schools, 5) Quidditch Team Members, 6) House Elf
Connecting Back
  • What do you like about this magical “thing” compared to Muggle social media?
  • What do you dislike about it?
  • Imagine using the magical “thing” we discussed earlier. How do you think it could resolve some of the issues people might face with Muggle social media?
  • How might this new magical “thing” better support your needs for self-presentation and identity management compared to traditional social media?
  • What challenges might arise in using such a “thing” at Hogwarts?
Thank you so much for your time and insights. Your contributions are incredibly valuable to our research. Do you have any final thoughts or questions before we conclude?
Appendix B Participant Interview Summary
Table 3. Key Features for Envisioned Social Media from Participant Interviews
PID Envisioned Social Media Features
P01 Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates, pinned messages, and casual interactions; Mailbox updates; Public areas for open interaction & Interest-based rooms for community engagement; Area-based access control; Music & mood reflection in rooms; News hub or enchanted map for real-time global updates.
P02 Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
P03 Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; Floating holographic interface; VR-like navigation; Selective visibility & privacy settings; Muggle guest system in selective areas with limited abilities; Shopping district & immersive commerce; Immersive media; Customizable avatars & dynamic appearance.
P04 Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods and activities; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
P05 Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
P06 Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers like knocking on a bookshelf or moving objects in a set sequence; House elf companion (AI assistant); Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
P07 Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
P08 Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
P09 Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Customizable bulletin board for updates; Moderation via magical guardians; Room of Requirement-inspired group chats.
P10 Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
P11 Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
P12 Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; 3D interactive messaging & calls; Interest-based connection zones; Engagement-based prioritization.
P13 Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
P14 Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Mutual unfollowing & privacy indicators; Authentic avatars.
P15 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
P16 Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
P17 Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
P18 Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
P19 Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
P20 Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
P21 Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking, teleportation, and environment transformation; Immersive social events; Space-based access control.
P22 Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
P23 Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; "Hall of Fame" memory posts.