### v2

# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

Abstract.
We investigate youth visions for ideal remote social interactions, drawing on co-design interviews with 23 participants (aged 15-24) experienced with 3D gaming environments. Using a Fictional Inquiry (FI) method set in the Harry Potter universe, this research reveals that young people desire social media that functions more like immersive, navigable shared social spaces. Across these interviews, participants identified six key priorities for meaningful social connection over social media: intuitive social navigation, shared collaborative experiences, communal environments fostering close relationships, flexible self-presentation, intentional engagement, and playful social mechanics. We introduce the spatial integrity framework, a set of four interrelated design principles: spatial presence, spatial composition, spatial configuration, and spatial depth. Together, these principles outline how online spaces can be designed to feel more like meaningful environments, spaces where relationships can grow through shared presence, movement, and intentional interaction. Participants also described the FI process itself as meaningful, not only for generating new ideas but for empowering them to imagine and shape the future of social media.
social media; fictional inquiry; youth; design
conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3ccs: Human-centered computing Collaborative and social computing
1. Introduction
Meaningful social connections play a critical role in young people’s development and well-being (Hawkley and Cacioppo, 2010), and many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n. d.]). However, recent trends indicate that social media platforms are increasingly becoming less focused on facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025). Initially designed around sharing personal updates with friends, these platforms now often emphasize entertainment, news, and engaging larger, less intimate audiences. Meta’s own data illustrates this shift: between 2023 and 2025, the share of time users spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025). Mark Zuckerberg described this change as aligned with the company’s focus on supporting users “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
This shift away from personal interactions raises concerns about the future of meaningful online relationships, particularly for younger users who have grown up with social media as a central part of their social lives. Most youth today engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023). Despite this high engagement, many young users report dissatisfaction with the quality of their social interactions, not because they wish to reduce their usage, but because they expect these platforms to better support their social connections (Kim et al., 2024b). Current platform designs frequently prioritize passive consumption of content over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
Given the central importance of social connections, particularly for youth, and recognizing social media’s vital role in facilitating these interactions, it is increasingly important to explore new design possibilities. Although young people rely heavily on social media, they frequently express dissatisfaction with their current experiences, indicating a significant gap between their social needs and the interactions enabled by existing platforms. Current mainstream social media platforms employ highly standardized design patterns, which limit our collective imagination and contribute to feelings of pessimism among both younger and older users. Consequently, there is a compelling need for innovative designs that complement existing platforms by placing a stronger emphasis on meaningful interpersonal connections. This prompts the critical question of understanding what types of social interactions current designs might be overlooking or inadequately supporting.
To address this question, we conducted a series of co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007). We involved 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n. d.]) and Roblox (Roblox Corporation, [n. d.]). Participants were asked to collaboratively imagine a social media platform specifically designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This imaginative context allowed participants to step away from the constraints of existing platforms and envision more varied and meaningful social interactions. Additionally, because 3D gaming environments are widely recognized as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated that participants could leverage insights from their gaming experiences to propose novel social media interactions. Specifically, our research aims to answer the following question: What types of interactions do youth familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
Our findings reveal that youth with 3D gaming experiences envision ideal social media platforms not as immersive, navigable environments designed to feel more like neighborhoods, community centers, and shared social spaces. Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming. Together, these visions highlight the youth’s desire for social media that feels more relational, meaningful, and personally rewarding.
We contribute empirical findings on what youth perceive as the most significant shortcomings of current social media platforms, as well as what they desire to experience on social media. Additionally, we propose the concept of spatial integrity for social interactions in online spaces, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. These dimensions include spatial presence (the ability to understand the space and identify who is present), spatial composition (the meaning and significance of the space), spatial configuration (the organizational structure of the space), and spatial depth (the extent of information that can be perceived within the space). Furthermore, we demonstrate that the FI method, in the context of social media design, not only prompted creative idea generation and enabled participants to transcend current social media design patterns but also empowered them to envision the positive potential of social media, leaving them feeling more hopeful about its future possibilities.
2. Related Work
2.1. Meaningful Social Connection in Online Environments
2.1.1. Defining Meaningful Social Interaction
Litt et al. (Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” While these interactions are described as “deeply subjective,” the primary distinction from non-meaningful ones lies in the presence of impact. This impact extends “beyond the immediate interaction itself,” typically involving positive emotions, the exchange of information, or concrete benefits.
The study identified three key types of impact: “emotional impact” (involving feelings, mood changes, or relationship strengthening), “informational impact” (relating to the exchange of knowledge, advice, or understanding), and “tangible impact” (providing concrete benefits such as goods or services). In contrast, non-meaningful interactions were often explicitly described as lacking any such impact.
Factors like the presence of strong ties (e.g., friends and family) and community ties, as well as participation in shared activities, were found to be associated with meaningfulness. However, these appeared to act more as facilitators that enable impact rather than as the primary source of meaning. Other attributes quantitatively linked to greater meaningfulness included whether interactions were planned in advance, memorialized (for example, through photos or videos), and synchronous. Interestingly, the study found that the communication medium itself, whether online or offline, was less important in predicting meaningfulness after accounting for these other characteristics.
While in this paper we treat meaningful social interaction as any interaction that participants say they would actively seek out in an ideal scenario. reflecting their subjective or perceived meaningfulness, we expected these themes to echo in how our participants talked about their experiences.
2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
Social media platforms are often seen as a primary means of connecting with others, especially for young people (Boyd, 2014a). However, despite the widespread use of these platforms, many users report feeling lonely or disconnected (Twenge et al., 2019; Primack et al., 2017). This disconnect highlights a significant gap between the intended purpose of social media and the actual experiences of users.
Social connection is just one of the many reasons people turn to social media, but it remains one of the most important. From staying in touch with close friends to meeting new people or maintaining community ties, users often come to these platforms seeking connection. Mainstream platforms like Instagram (All, [n. d.]), YouTube (you, [n. d.]), and Discord (discord page, [n. d.]) support some of these needs, but they are designed to address a wide range of user motivations, including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022). From a uses and gratifications perspective (Katz et al., 1973), even within a single platform, users pursue diverse goals, ranging from building social capital (Phua et al., 2017) to engaging in general media consumption (Menon, 2022). As such, although people may approach these platforms with the intention of connecting socially, they may encounter experiences that prioritize other forms of engagement, potentially leaving their social needs unmet.
This gap highlights the importance of developing a clearer understanding of what it means to support meaningful social connection in online environments. While social media offers both risks, such as social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b), and benefits, such as support for marginalized groups (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and opportunities for identity development and self-expression (Davis, 2012; Lee et al., 2022), social connection remains a central motivation for many users. Particularly for young people, who increasingly rely on digital platforms for social interaction, it is critical to move beyond treating connection as one of many competing platform goals and toward a more focused understanding of how platforms can meaningfully support users’ social relationships.
2.2. Social Technology Design Explorations in HCI
2.2.1. Designing for Meaningful Social Connections
HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features. Social media platforms such as BeReal (Kim et al., 2024b), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression. This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction. However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
At the feature level, efforts to support meaningful communication include designs that encourage effortful communication (Liu et al., 2021; Zhang et al., 2022) and those that foster feelings of genuine connections (Stepanova et al., 2022). While these features are important, social media experiences are often holistic, involving a complex interplay of multiple design aspects and user behaviors. As a result, such features alone cannot fundamentally alter the overarching narratives or perspectives surrounding social media.
Beyond archetypal social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions. Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments. Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions. However, these platforms often are not generally recognized as social media in the generic sense.
What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear, especially when considering the possibilities offered by emerging technologies such as spatial computing platforms. Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
2.2.2. Video Games as a Source of Meaningful Social Connection
Video games are increasingly recognized as valuable environments for fostering meaningful social connection, particularly among youth. Large-scale survey research shows that video games serve as a major venue for the development and maintenance of friendships, especially for boys (Lenhart, 2015c). Teens not only play games with friends they know in person but also frequently make new friends through online gameplay (Lenhart, 2015a), with more than half of teens reporting they have met at least one friend online (Lenhart, 2015a, b). These interactions extend beyond gameplay itself, supporting ongoing social relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
Research also highlights the social and emotional benefits of gaming, including opportunities for cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014). Certain game designs promote prosocial outcomes by encouraging players to help and support others, both in-game and beyond (Gentile et al., 2009). Moreover, games are not limited to surface-level entertainment; they can also offer deeper, eudaimonic experiences that involve meaning-making, personal growth, and feelings of connection with others (Daneels et al., 2023). These findings suggest that video games can be powerful tools for fostering meaningful social connections, particularly among youth, and highlight the potential for game-like features to enhance social media experiences.
2.3. Recognizing and Overcoming Social Media Design Fixation
Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design. While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking. This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
Social media design is no exception (Sundaram, [n. d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020). Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps (Oth, [n. d.]). However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024b) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)). These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
2.3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios. These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings. The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints. While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
According to Dindler and Iversen (Dindler and Iversen, 2007), implementing this method involves three key steps: first, clearly defining the inquiry’s purpose, whether it’s for staging design situations, exploring future ideas, or driving organizational change; second, developing an appropriate narrative that participants are comfortable with while being distinct enough from current practices to encourage new thinking; and third, creating a compelling plot that introduces tension or conflict within the narrative to motivate specific workshop activities. Unlike traditional role-playing approaches, FI encourages participants to maintain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities while drawing from their personal motivations and capabilities, all while circumventing typical societal constraints.
Further, one of the most powerful ways to overcome design fixation is analogical transfer (Novick, 1988; Buchanan, 1992). We see video games as a rich source of analogies for social media design, offering unique affordances and interaction patterns that can inspire new approaches to fostering meaningful social connections. By examining how video games facilitate social interactions, we can identify design patterns that can be adapted to social media contexts, helping to break free from design fixation and explore novel alternatives.
3. Method
We conducted co-design interviews with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms. All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n. d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers. During the interviews, we used the FI method to scaffold participants in generating design ideas.
3.1. Material: The Narrative Framework
Figure 1. Miro board mapping real-life relationships to Hogwarts-themed groups, designed to immerse participants in the Hogwarts-inspired FI narrative.
[…]
This is a visual chart designed to organize relationships and connections in the Harry Potter universe. There are three columns of labeled sections. In the middle, there are four sections, each representing one of the Hogwarts houses: Gryffindor (red), Hufflepuff (yellow), Slytherin (green), and Ravenclaw (blue). Each house box has its crest. To the left, there are two smaller boxes. The top-left box is labeled “Best Friends / Romantic Relationship(s),” featuring small images of Harry Potter, Hermione Granger, and Ron Weasley, with a space for a name or nickname. Below this, there is a smaller box labeled “House Elf,” featuring a picture of Dobby holding a glowing object and an area for a name. To the right, there are two more boxes. The top-right box is labeld “Family,” containing images of Sirius Black, James and Lily Potter, and the Dursley family. Below this, another box is labeld “Profs & Other Adults,” displaying images of Severus Snape, Dumbledore, Professor McGonagall, and Hagrid. At the bottom, there are three large sections. The leftmost section is labeled “Muggles” with a blank space. The middle section, labeld “Students from Other magical Schools,” displays crests from Beauxbatons, Durmstrang, Ilvermorny, and Castelobruxo. The rightmost section, labeled “Quidditch Team Members,” features an image of a Quidditch team in red robes. Each section contains an empty space for names or details to be added.
For our study, we selected the Harry Potter world as the narrative framework for our FI method. The Harry Potter series follows the journey of an orphaned boy who discovers he is a wizard and, as a result, is invited to attend “Hogwarts School of Witchcraft and Wizardry,” where he forms deep friendships and gradually uncovers his destiny as he clashes with the story’s antagonist.
Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI interviews. The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement. The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells. Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies. Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building. The characters’ experiences or life transitions, which parallel real-world moves between schools or locations, allow for exploring how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
3.2. Procedure
The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI method. The interviews averaged around 90 minutes. We began the interview by asking which 3D games they had played and what they liked about those experiences. We also asked questions about Harry Potter, such as their favorite book in the series. This served to break the ice and help them start immersing themselves in the Harry Potter world.
Following this introduction, we presented the FI narrative: participants were asked to envision ways to support social connection within the Hogwarts world as a non-Wizard (i.e., “Muggle” (mug, [n. d.])) trying to stay connected with everyone across Hogwarts and the Muggle world. We then shared a Miro (mir, [n. d.]) board (Figure 1) where participants could sort their friends and family into categories inspired by the Harry Potter series, such as fictional houses (e.g., Gryffindor, Hufflepuff), teams (e.g., Quiddich Team members), and old friends (e.g., muggles), reflecting fictional houses, teams, and communities from the story. This activity encouraged them to think of their relationships within the Harry Potter context and was intended to further enhance immersion in the fictional design space. Next, we asked general questions about their experiences with social media, focusing on what they found meaningful or meaningless and their vision for an ideal social media platform.
During the design phase, participants were prompted to describe the capabilities their ideal magical social media would have and what aspects of their lives they would choose to share. Throughout this phase, we sought to keep participants immersed in the narrative, referring to existing platforms as “Muggle social media” and maintaining consistent engagement with the Hogwarts setting, as outlined by the fictional inquiry guidelines. Finally, we concluded by asking participants to reflect on how the social media they envisioned for Hogwarts compared to Muggle social media. The semi-structured interview protocol is available in Appendix A.
3.3. Participants and Recruitment
We used a combination of purposive and convenience sampling to recruit participants for our study. The first and second authors shared the recruitment materials (that included a link to a screener survey) on their social media accounts, while the first author also distributed flyers across our institution’s campus.
Participants were required to be between 15 and 24 years old, in accordance with UN (United Nations, [n. d.]) and WHO (who, [n. d.]) definitions for youth, and to consent to be video or audio recorded during a 90-minute interview. Video recording was requested following the FI method guidelines, enabling us to analyze gestures if necessary. We included both teenage and young adult participants (ages 15 to 24) to capture perspectives from distinct life stages, as these groups often navigate significant transitions and share a focus on identity exploration, self-presentation, and relationships.
They also had to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox). While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
Table 1. Demographics Summary. Abbreviations: Familiarity with Harry Potter - Expert (“I remember every detail about the plot, characters, and magical aspects (like how the castle moves, spells, magical creatures, etc.).”), Proficient (“I remember most details …”), Familiar (“I remember some details …”). Social Media Platforms – IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace). Reasons for Using Social Media – Connect (Connecting with friends and family), Network (Networking for professional purposes), News (Staying updated with news and current events), Share (Sharing content (photos, videos, etc.)), Community (Joining communities or groups with shared interests).
PID Familiarity with HP Social Media Platforms Reasons for Using Social Media Gender Age Ethnicity
P01 Proficient IG, DC, BR, RD Connection, Network, Entertainment, Share Woman 19 Asian or Asian-American
P02 Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI Connection, Network, Entertainment, Share Woman 20 Asian or Asian-American
P03 Proficient IG, DC, TT, X, RD, TH, NP Connection, News, Entertainment, Share, Community Woman 18 Black or African-American
P04 Expert IG, DC, SC, TT, BR, X, RD Connection, Entertainment Woman 16 Asian or Asian-American
P05 Proficient IG, DC, BR, RD, FB, TH Connection, News, Entertainment, Community Woman 18 White
P06 Familiar IG, DC, SC, TT, RD, FB Connection, News, Entertainment, Community Non-binary 18 White
P07 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, News, Entertainment, Share Woman 17 Hispanic or Latino, White
P08 Proficient IG, DC, SC, RD, FB Connection, Entertainment, Share, Community Woman 19 White
P09 Proficient IG, DC, TT Connection, Network, Community Woman 21 Asian or Asian-American, White
P10 Proficient IG, DC, SC, TT, BR, RD, TH Connection, News, Entertainment, Share, Community Non-binary 17 Hispanic or Latino, White
P11 Familiar IG, DC, TT, X, RD, FB, TH Connection, News, Entertainment, Share Woman, Non-binary 19 Native American or Other Pacific Islander, White
P12 Familiar IG, DC, SC, TT, X Connection, Network, News, Entertainment, Share Man 19 Asian or Asian-American
P13 Familiar IG, DC, SC, TT, BR, X, RD, FB, TH Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P14 Expert IG, DC, SC, TT, BR, X, FB Connection, News, Entertainment Woman 18 Hispanic or Latino
P15 Proficient IG, DC, SC, BR, X, RD Network, News, Share, Community Woman 20 Asian or Asian-American
P16 Familiar IG, DC, SC, BR, X, RD Connection Man 18 Asian or Asian-American
P17 Expert IG, DC, SC, TT, X, RD, TH Connection, News, Entertainment, Share Woman, Non-binary 17 Hispanic or Latino, White
P18 Familiar IG, DC, SC, TT, X, FB Connection, News, Entertainment Woman 23 Asian or Asian-American
P19 Expert IG, DC, SC, TT, X, RD, FB Connection, Network, News, Entertainment, Share, Community Self-described, GenderFluid/Trans Man 19 White
P20 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, Share, Community Woman 17 Black or African-American
P21 Expert IG, DC, SC, RD Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P22 Familiar IG, DC, SC, TT, X, RD Connection, News, Entertainment, Share, Community Woman 19 White
P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
A total of 84 individuals completed the screener survey. We reached out to the 48 respondents who met the recruitment criteria, and 23 participants provided consent and completed the interview procedure. The mean age of the participant group was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) identifying as women, 3 (13.0%) as men, 2 (8.7%) as non-binary, and 3 (13.0%) that were gender fluid. The full list of participant information is available in Table 1.
3.4. Data Analysis
We analyzed the interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides a systematic yet adaptable framework for interpreting qualitative data. The method’s independence from pre-existing theoretical frameworks made it particularly suitable for our study, given our study’s exploratory focus on youth-envisioned social media design through the FI method. It enabled us to effectively capture and analyze the diverse perspectives shared by our participants.
Our analysis began with the first two authors independently coding the same two transcripts line by line. During this phase, the codes were descriptive and closely adhered to the data. The authors convened to discuss and resolve any discrepancies in their coding. Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024). After four coding iterations, we reached saturation in our codes. Using this final set of codes, the first author proceeded to code all the transcripts. All versions of the codes and the final codebook are available in the supplementary materials.
3.5. Ethical Considerations
Given J.K. Rowling’s controversial statements regarding transgender individuals (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in our recruitment material and screener survey and reiterated our position during the interview procedure.
4. Results
Informed by their experiences with 3D games and familiar media worlds, our participants imagined forms of social interaction that extend beyond what is typically supported on mainstream platforms like Instagram. Through the speculative frame of Social Media at Hogwarts (SMH), they explored how digital environments might enable more embodied, relational, and emotionally meaningful forms of connection. Their visions converged around six interrelated themes, which, collectively, contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs, ultimately expanding their perception of the horizons for social connection on such platforms.
4.1. Intuitive Social Navigation
Youth described seeking interactions that feel natural, embodied, and socially meaningful, reflecting the way people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape how they connect with others, express themselves, and manage privacy in intuitive and socially familiar ways.
4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Participants described how embodied, immersive qualities such as realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space that fosters presence, emotional connection, and socially meaningful interaction. In contrast, current social media (which we will refer to as “Muggle Social Media” or MSM for short), was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 similarly noted that MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” P03 described SMH as “more immersive and more awesome” than MSM, which they saw as inherently limited “because it’s 2D.” P16 emphasized that meaningful social connection requires “talk[ing] to you and see[ing] your face,” something they believed SMH could make possible.
For others, the fulfillment came from the embodied experience of physical actions. P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They also added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 similarly reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” As P06 put it, SMH “feels more real” because you “walk around a physical world” and “feel like you’re actually in the area of these people.”
This sense of presence was felt even in seemingly passive actions. As P01 explained,
“I don’t know, there’s just a feeling about it that makes it different. I can’t really explain. Depending on 2D and 3D, but I feel like in 2D, you can’t really change anything about their homepage. You just observe it. The most you can do is add a like or something, and that would change their like count. But then they would know who you are, so that takes away the anonymity. But if you’re just looking at it, then you’re not really affecting anything. That’s just for your knowledge. But in a 3D setting, even if you don’t leave any physical footprints, in a way, they’re still there because you opened the door, you went in, you sat down. Maybe just that mark in a way that you’ve left on the place takes away from the anonymity. I don’t know if that makes sense but.” (P01)
For P01, 3D interaction not only supported more embodied engagement but also introduced a felt presence that lingered, even when no direct interaction took place.
4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed that using spatial and visual cues modeled after real-world environments could make managing access and privacy feel more intuitive, socially natural, and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
“I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home? If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
While managing visibility across multiple rooms might seem more complex, participants believed this approach would feel intuitive and natural. P11 suggested that spatial contexts could lower the efforts required for privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.2. Shared Experiences
Participants emphasized that shared experiences, whether through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
4.2.1. Fostering Shared Memories Through Collaborative Activities
Participants described shared activities as a meaningful foundation for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. For example, while playing a simple emoji game was described as “just something fun we can both do,” it lacked a true shared experience since “we don’t actually see each other’s screens” and would just “screenshot our scores and send them to each other.” By contrast, they described Roblox as a space that fosters shared moments that become lasting memories, like “do you remember that time when that kid in the voice chat said this,” or “when we got scared by that jump scare.” They explained, “it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
Similarly, P12 described the value of collaborative projects, such as “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting it with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than simply “posting” content individually.
The value of shared experiences and memories was not limited to completing tasks or reaching goals. Even simple activities like walking together were seen as meaningful opportunities for connection. P03 imagined moments where you could just “walk with your friend,” as if taking a “nature walk” (P14). P21 echoed this, describing how walking itself could spark memorable conversations and shared moments: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Participants also saw low-pressure, ambient forms of social interaction, such as co-existing in shared spaces or listening to music together, as natural and organic ways to build connections and shared memories. P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background. They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations. P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.” Even “mindless scrolling” became meaningful when shared. P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music would often facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They came up with the idea reflecting on their own experience:
“At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music. And it was one of the most awesome nights because we all knew in that moment that collectively we all enjoyed what we were hearing and it was just good. I don’t know how else to describe it. It was just nice.” (P19)
They added that music could also serve as a “conversation starter” while also providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
4.3. Communal Ambience
SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships. Participants described several design features, such as neighborhood structures, shared gathering spaces, and stronger boundary controls, that could help cultivate this kind of communal, low-pressure environment.
4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to naturally foster social interaction and relationship-building. P22 described in detail:
“So, in a virtual world, you might have your house, and then you could go to a community board or town square with postings for different groups … From there, you could travel to those group spaces. I don’t know if you’ve ever played or seen Club Penguin, but in that world, you had your igloo to decorate and could visit the town … The town had different buildings for different activities. I’m imagining something like that, where you have your house where people can visit and talk with you, and if you want to meet new people, you can leave your house and go out into the town.” (P22)
Youth envisioned SMH as also reflecting the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties. Specifically, some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction. One participant elaborated:
“And in this magical world, I imagine the people around you in your neighborhood would be from the same house as you, kind of like how people live together in dorms on a college campus. The nearby stores would be the ones your group usually goes to. So you’d share that sense of togetherness because you all go to the same stores or belong to the same house. I think that would help build a real sense of community.” (P03)
Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them. This design approach of aligning physical landscapes with emotional closeness was seen as a way to streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
Participants imagined that in addition to private homes and personal spaces, SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating opportunities for connection. P05 described these areas as “neutral spaces”, not tied to any specific activity or relationship status, where groups could gather for a range of purposes. They imagined “some places to sit, and some kind of flat table-like surface” that could support casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces, they explained, would function as “a third, neutral territory” where people could interact more organically, without the friction of managing invitations or switching platforms.
Participants described shared social spaces within neighborhoods, such as common rooms, courtyards, or clubhouses, where residents could casually interact outside their personal spaces. P01 described leaving one’s home to gather in a “courtyard or game room” at the center of the neighborhood. Drawing on examples from Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms that cater to different interests, such as a Quidditch field or library, creating opportunities for people to meet others through shared activities. P09 similarly imagined a “virtual dining room” that would allow people to casually talk while doing something familiar, like eating together. Unlike the more segmented house-based accommodation at Hogwarts, this dining space was described as open and inclusive, allowing people to connect across social groups.
Others described third places located in public areas that supported casual, serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look! This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized the importance of filtering out irrelevant content, such as celebrity drama or political commentary, and prioritizing updates from close friends and trusted communities. P19 criticized how platforms like Instagram have shifted away from their initial focus:
“We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page. We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
Several participants advocated minimizing celebrity influence. P02 stated, “I don’t want celebrity stuff… I don’t want any celebrity drama,” while P01 suggested removing celebrities altogether if the platform prioritized personal enjoyment. If included, they proposed isolating them to a “celebrity zone.” Prioritizing posts from close friends was a key preference. P07 explained, “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH offering greater control over interactions and content, noting it would be “more designed for people you actually know… like your friends, classmates, or Quidditch team,” rather than celebrities or acquaintances. P19, similarly, envisioned that on SMH, they could “filter out for friends and family, and even then condense your circle a little bit more.”
4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Participants described how themed, socially grounded conversational spaces designed for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
“Anyone you’re talking to in general, just emphasizing conversation between individuals more, rather than just posting on a thread. Like how I talked last time about Reddit, you know, there’s a thread, and then you make a comment on it, but there are a hundred other comments, so no one’s going to respond to you. I want to focus more on having actual conversations with people, making real connections, and more individualized chatting. I find it’s way easier to get to know people a lot better.” (P22)
Participants described how the design and feel of conversational spaces could make social interactions on SMH feel safer, more personal, and more inviting than on existing platforms. Many valued the idea of walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to a broad, undefined audience. P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations. They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
Participants also imagined personalizing these conversational spaces to make them feel more emotionally connected. P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories,” making even casual conversations feel more intimate. P09 expanded on this idea, imagining a “Room of Requirement”-like system (roo, [n. d.]) where users could create custom rooms on demand to match the “purpose” of their conversation. Participants also valued interest-based spaces, similar to Discord (discord page, [n. d.]), where a “common point” (P13) or “good talking point” (P07) could help break the ice and give conversations direction.
4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
To ensure SMH feels like a safe, personal environment, participants stressed the need for more comprehensive boundary controls beyond privacy tools found on existing platforms that they perceived as limited and inconsistent. P22, for example, explained that even with contact syncing turned off on Instagram, the platform still suggested their Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they were trying to avoid, based on mutual follow relationships. They stressed the need for stronger, more thorough blocking controls, including the ability to block “alternate accounts” and fully remove any trace of a blocked person’s presence, explaining, “I don’t wanna be able to tell that the person I blocked is even around me.” P14 similarly pointed out that even private accounts on Instagram still expose personal content, like Story Highlights, to “a random person”, revealing perceived gaps in platform-level privacy protections (Bojkov, 2019).
Participants expressed a desire for stronger control over their own visibility, imagining features that would let them opt out of being seen or discovered entirely. P03 suggested a feature like an “invisibility cloak” for moments when users “don’t want anyone to see you.” P07 proposed a “super private” profile mode where users “can’t even be searched… like you’re not even on the app” unless they choose to make themselves visible.
Preventing unsolicited contact from strangers was another recurring priority. P22 proposed a model inspired by Discord, where only people you’ve already interacted with can contact you, and others “can’t even be discovered unless you have something in common.” This idea resonated with P13, who argued that new connections should require mutual consent, suggesting that people “should only be able to add you if you meet in person and exchange a QR code.” They shared a personal experience of being able to find a stranger online based on just their first name, reflecting that “it’s too easy to find people on social media.”
4.4. Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting the full range of who they are, without the pressures of curation or performance.
4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Participants consistently described customizable avatars and personal spaces as meaningful tools for self-expression, offering more creativity, personality, and authenticity than the static profiles or limited settings found on MSM. Unlike MSM, which typically offers only “two modes” (P10), i.e., “light and dark” (P10), participants imagined a much richer and more immersive experience. For example, P21 envisioned a fully customizable world where users could build their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” These spaces could be shared with friends as a way of “showing a little part of yourself.” Similarly, P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture,” to reflect their personal interests.
Customizable “3D avatar[s]” (P11) and profiles were also seen as key ways of showing individuality. P22 reflected on the early days of Myspace (mys, [n. d.]), describing how they enjoyed coding their profile layouts and changing “the font and really make it reflect me.” Even small details, like choosing a username, were described as ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted that such forms of expression are often “mainly for you,” not just for public display. P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right? So I think that’s part of you.”
Beyond visual customization, participants imagined SMH supporting more ambient, indirect forms of expression, offering subtle cues about what someone is doing, feeling, or thinking without needing to make a formal post. For instance, P01 described wanting their avatar to reflect real-time states, like the outfit they are wearing or the song they are listening to, so that friends could intuitively notice things like, “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ Similarly, P10 imagined that friends could pick up on changes in their interests simply by observing how they update their virtual space over time, providing “an insight into… what I’m thinking.”
Several participants highlighted the contrast between the flat, compressed feel of 2D platforms and the immersive qualities of 3D environments. P10 noted that in 3D, others can see “as much stuff as you can see in 3D,” making it feel richer and more complete. P16 added that VR creates a sense of presence that feels more like meeting “real people… compared to the profile pictures,” because you can pick up on their “mannerisms” and notice details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
Participants also described how these kinds of choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests. As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
Drawing parallels to the physical world, P13 pointed out how personal spaces like bedrooms carry a sense of intimacy and authenticity that translates well to SMH. They explained:
“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate. It’s where you spend most of your time, and it shows your personality the most. It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes or 30 seconds the way you can on Instagram. If you wanted to add someone there, all you’d have to do is remove a couple of things. It’s easier than physically changing a space.” (P13)
Finally, participants appreciated these ambient and low-pressure forms of expression for their lower social stakes. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling like they are performing for attention. They described this kind of expression as “more justified” and less “embarrassing” than posting content just to be seen by others.
4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides of themselves across various spaces and relationships, and reducing the pressure to maintain a single, polished identity. Rather than being limited to a single, polished identity, as is often the case on MSM, participants wanted ways to represent the full range of their interests, identities, and relationships. P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are. P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
4.4.3. Sharing Everyday Activities and Lived Experiences to Deepend Understanding and Connection
Participants valued the ability to share casual, everyday moments and rich, embodied memories, filling the gaps between the highlights usually shared on MSM into a more natural, meaningful extension of daily life and personal connection. P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants also imagined SMH supporting the sharing of more personal, embodied memories as a way to deepen understanding and reduce misinterpretation. P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n. d.]) as a way to create stronger emotional connections. Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully. P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
Some participants imagined that these shared memories could include sensory or emotional details, not just visuals. P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication. Others compared it to live streaming personal experiences. P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record. P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Participants emphasized that flexible, contextual privacy is essential for supporting authentic self-presentation, allowing people to manage who sees what, when, and how, without feeling overexposed or misrepresented. They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life. P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
Participants also reflected on the challenge of managing different audiences and social contexts on MSM and imagined SMH supporting more seamless transitions between identities and spaces. P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts. Building on this, participants imagined contextual identity management, where users could shift how they present themselves based on the audience or situation. P03 proposed a model where everyone maintains a “base” identity but customizes additional layers. P22 similarly described having different avatars or profiles for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
Participants also stressed that contextual privacy should extend beyond people-based access to include temporal and relational factors. P14 proposed limiting access to different stages of one’s past identity, saying, “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” Others, like P21, suggested gradual openness, where people could selectively reveal more of themselves over time. As they described, “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.”
Participants also highlighted the importance of individual control over these transitions and boundaries. P14 explained that privacy settings should not be one-size-fits-all, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy as not just a technical feature, but as a core requirement for supporting the kinds of personal, meaningful, and low-pressure interactions youth hoped SMH would make possible.
4.5. Intentional Engagement
Participants described their growing frustration with how current social media platforms prioritize passive consumption over meaningful connection. In response, they envisioned SMH as a space that promotes intentional, socially rewarding interactions. This section explores four key themes that reflect their desire to refocus social media on authentic connection, purposeful engagement, and meaningful use of time.
4.5.1. Encouraging Deliberate Navigation and Content Discovery
Participants imagined SMH as a space that encourages deliberate, effortful discovery, replacing the passive, algorithm-driven feeds of MSM with experiences that require users to actively choose what and who they engage with. P22 described their frustration with Instagram, explaining that “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage more intentional, purposeful engagement, both when browsing content and when navigating the virtual world.
P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful. Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
Participants also imagined features that would prompt users to actively choose what they wanted to engage with before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was also seen as key to this intentionality. Participants valued having to walk, explore, and navigate to find people, spaces, or activities, turning even small actions into social moments of presence and care. P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling.
This desire for intentional interaction extended to how users move through the platform itself. Participants imagined physical navigation, such as walking or traveling through virtual space, as a way to make engagement feel more meaningful. P02 explained that they would prefer slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” While features like “spells” (P06) or the “Marauder’s Map” (P21) were imagined as convenient options, participants valued the ability to walk and explore, turning simple actions like visiting a friend or browsing a shop into shared social moments. P03 explained that this kind of embodied interaction feels “more connected than regular social media” (P03) because it requires active presence rather than passive scrolling.
P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them. They described SMH as a world where you might “walk through a town to get there,” explaining, “It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
Participants believed that this kind of effortful navigation would naturally discourage casual lurking or influencer-driven engagement, shifting the platform’s focus toward genuine connection. As P05 put it, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond how users navigate content, participants also emphasized the importance of shifting the platform’s core purpose from endless consumption toward meaningful social connection. Participants expressed frustration with how current platforms promote mindless, passive consumption, and described their hopes for SMH to instead foster more meaningful, intentional connections through active social engagement. They described frustration with how easily time slips away on platforms like Instagram and TikTok without leaving them feeling unfulfilled. P10 noted that social media often feels “meaningless” when they “didn’t learn something,” and P18 reflected on the regret of “mindlessly scrolling.” P14 similarly described spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
A major contributor to this feeling was the infinite scrolling design of MSM. P11 explained that “there’s always more content on social media,” which fuels the fear of missing out and the urge to keep scrolling. P14 critiqued TikTok for exposing users to “clickbaity” and often low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
In contrast, participants described active, social engagement, even if casual, as far more rewarding. P16 shared that “anytime not scrolling” but instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this, noting that meaningful use comes when they are “actually talking to someone,” especially to stay connected with people “not in close proximity.” They reflected that while users might spend more time on a platform like SMH, that time would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to passively watching “50-second TikTok videos of somebody making slime for like an hour.”
Participants also contrasted passive consumption with active participation in interactive environments. P10 explained that while doomscrolling feels unproductive, activities like playing Minecraft feel different because they offer “progress” that can be “saved and come back to.” P06 echoed this sentiment, noting that they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
Reflecting on SMH’s potential to create this kind of active, engaging experience, P09 acknowledged that “time spent might not be mitigated,” but emphasized that “the way you’re spending it” would make the difference. They explained that SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting that meaningful engagement, rather than passive consumption, was what ultimately mattered for them.
4.6. Playfulness and Whimsy
Participants described SMH as a platform that feels youthful, playful, and deeply immersive, offering not just entertainment but opportunities for creative self-expression, social bonding, and emotional well-being. By embracing whimsy and delight, SMH was seen as making social media feel more personal, less pressured, and more meaningful.
4.6.1. Embracing Creative Joy and Playful Self-Expression
Participants imagined SMH as a platform that transforms social media into a space for creative joy and playful self-expression, offering users the freedom to experiment with identity, explore personal interests, and connect with others through creativity. Several participants described how SMH offered far more than static profiles or simple feeds. P06 reflected that “everyone likes sharing parts of themselves,” appreciating how the platform allows users to shape not just what they say, but how they appear and interact with others. P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
For some, this act of personal customization felt like a meaningful process of identity exploration. P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression. P22 added that SMH offered something they hadn’t found anywhere else, explaining, “I can really be creative here… I can make something that I haven’t made anywhere else, you know. That’s where the fun aspect comes in, really, for me.”
This spirit of creativity extends beyond the individual to the wider environment and community. Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa! They built that, that’s so cool” (P06). SMH felt alive and engaging whether users were socializing or spending time alone: “it’s not just like, Oh, the group chats are only fun when there’s other people here. No, it’s fun when we’re by ourselves, too, we get to do our own things” (P06). The interactive environment itself, where “you’re not just sitting there chatting, you’re also walking around the space… you’re interacting with whatever else is in that area” (P06), offered more to do than conventional platforms. Whether through building, exploring, or engaging with shared content, SMH transformed social interaction into something playful, creative, and deeply engaging.
4.6.2. Designing Engaging and Playful Privacy Mechanisms
Participants imagined privacy not as a tedious and confusing task but as part of the playful experience, envisioning interactive, game-like mechanisms that turn boundary-setting into something fun, discoverable, and socially meaningful. P11 envisioned systems where users could control access to spaces through gestures like a “secret knock” or hidden objects that revealed “secret passages” accessible only to trusted friends. Others suggested knowledge-based gates, such as “secret handshakes” (P21) shared among friends. Drawing inspiration from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces only when correctly summoned.
These features were described not just as functional but as adding delight and novelty to social interaction. P05 compared them to “little Easter eggs,” remarking that “there’s that little kid part of everyone” that makes such features enjoyable. They contrasted these with conventional privacy settings on mainstream social media, which they described as “a wall of settings” that often feel confusing or uninviting.
4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities. P01 questioned the need for phones at all, asking, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting “holographic screens, like Jarvis in Iron Man,” or activating “panels” and “big screens” through spells. Others suggested alternative magical devices. P15 preferred VR glasses, describing them as “more natural,” “easier,” and “more fun.” Participants also imagined objects like a “makeup mirror” that could serve as a portal, a “watch” that immerses users into SMH, or a “mirror glass” for video calls or summoning friends with spells like accio.
These magical affordances were not only imagined as novel interfaces but also as ways to make extraordinary experiences more accessible. P09 highlighted that SMH could give people “experiencing stress” or lacking access to “really, really beautiful landscape and environments” the chance to walk around and “admir[e]” calming, scenic spaces. P16 imagined that users could take part in activities they might never otherwise experience, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
4.7. Expanding Horizons
Given the speculative nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what interactions on social media platforms could embody. Therefore, while they acknowledged potential challenges, such as the platform being over-immersive or distracting, they overwhelmingly viewed these downsides as worthwhile trade-offs for its many benefits. Overall, participants described SMH as an inspiring alternative or expansion of existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional. They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
4.7.1. Breaking Free from Platform Monotony with Novel Interactions
Participants saw SMH as a platform that offers better user experiences, departs from the repetitive structures of existing platforms, and inspires hope for what social media can become. For example, P06 shared how SMH combines several desirable aspects of social media in a cohesive way, creating an environment focused on meaningful interactions: “You get friends who you can talk to, keep up with, and stay connected to. And you get things you can customize. You kind of get the whole package there. I kind of like all of those here.”
Participants also appreciated how SMH offered a break from the monotony of existing platforms. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram. P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “novel,” like a “breath of fresh air.”
For some participants, SMH’s novelty lay in giving themselves permission to prioritize enjoyment over seriousness. P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun. For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
“Yeah, [I would prefer the 3D version] actually 100%. I feel like,… the only reason I said 2D is because I felt like the need to be like a mature adult. But who cares? I feel like something that’s more enjoyable would be [the SMH].” (P14)
P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms. P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness. P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
Through the FI process, participants realized they were not limited by the structures of existing platforms. Instead, they felt empowered to imagine social technologies that center joy, creativity, and authentic human connection, qualities they had once believed were beyond the reach of digital spaces. P10 pointed out that overcoming the design fixation around social media is crucial:
“The way that we can make social media better is if we expand the realm of what the definition of social media is. Instead of it being where someone just shares stuff, it can also be places where people connect, and people do this and that.” (P10)
Several participants noted that the FI workshop played a pivotal role in helping them gain a broader vision of social technologies and envision new possibilities for social media. P15 reflected on how their perspective evolved throughout the process:
“When we started, I literally was giving all these suggestions, which was completely the same thing which existed… I was repeating their concepts. So I think throughout this entire interview, I evolved my idea… actually thinking about what is new… So I think this has been a super hit. I like the entire conversation.” (P15)
They appreciated that SMH was not merely a “clone of the same thing… without actually adding anything else” but rather a concept that was “completely different… not even the same.” Similarly, P14 shared how the workshop shifted their initial skepticism:
“I don’t know what else you could do, because it’s online… so I really don’t know what else there could be to make it more meaningful.” (P14)
After engaging with the workshop, they felt empowered to imagine broader possibilities:
“I’m not really a tech person. So I don’t really think about [how to make social media better] that much. But I feel like this fake scenario… [helped] remind myself, ‘Girl, you have full control, this is no restrictions.’ After I got into that I… have a way bigger vision of what social media can become. (P14)
5. Discussion
Figure 2. 3D Gamer Youth’s Vision of Ideal Social Media Interactions
[…]
A Venn diagram comparing two categories of online platforms. The left circle represents platforms focused on social networking and broadcasting, with icons for Instagram, Facebook, and TikTok. The right circle represents platforms where youth engage in meaningful connections, with icons for Zoom, Minecraft, Discord, and Roblox. In the overlapping center, labeled “Social Media at Hogwarts (SMH)”, the diagram highlights youth’s vision of an ideal platform that combines the social visibility of networking platforms with the depth of meaningful connection found in gaming and community spaces.
Figure 2 captures how participants described their vision for ideal social media as something that pulls benefits from both types of platforms they already use. On one side are social networking and broadcasting platforms like Instagram, Facebook, and TikTok, which make it easy to stay socially visible but often feel shallow or performative. On the other side are platforms like FaceTime, Minecraft, Discord, and Roblox, where youth experience more organic and engaging connections, but where social interaction is often secondary to other goals like gaming or productivity. Our participants imagined “Social Media at Hogwarts (SMH)” as a platform that blends these strengths, creating a space where meaningful social connection effectively surfaces not just a peripheral effect, but a key focus of the platform.
5.1. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings build on prior research showing that while youth are often dissatisfied with archetypal social media platforms, their goal is not to reduce time spent connecting online. Instead, they seek spaces that support deeper, more intentional social interactions. Participants described social connections that feel genuine, fulfilling, and emotionally present and not performative. The changes in social media platform design that they sought were as follows:
  1. (1)
    Immersive Experiences Through Spatial Elements. Youth expressed a strong desire to move beyond “flat” social media toward spatially immersive experiences. They imagined platforms where relationships could unfold through presence in shared 3D environments, not just posts on a feed. They described the value of feeling physically co-present with others, such as walking together in a virtual world or talking in a space that feels alive and embodied. They also envisioned personal spaces and community-centered landscapes, such as personal homes or neighborhoods that mirror real-world social structures. Several imagined futuristic forms of presence, such as those of holographic displays or VR interfaces that make interacting feel natural and embodied, reducing the sense of separation created by screens.
  2. (2)
    Organic and Intentional Social Interactions. Participants described meaningful connections as unfolding through casual, low-pressure, and community-oriented interactions. They valued flexible and safe spaces for direct conversations, where people could gather in protected and themed rooms to get to know each other more personally. They also emphasized the importance of community gathering places and shared activity spaces, such as areas to play games, collaborate on projects, or simply hang out. Visions of these spaces encompassed both low-intensity ambient co-presence, such as sitting together quietly while listening to music, and public third places that enable serendipitous encounters with new people. Participants also described the value of effortful and intentional navigation, preferring to move through spaces purposefully rather than having content or people pushed to them by algorithms.
  3. (3)
    Expressive and Lower-Stakes Sharing. Participants sought ways to express their identities naturally and casually, without the pressure of curating formal posts. They valued personal area and avatar customization as ways to reflect their personality through how they decorate spaces or design their avatars. They also envisioned ways to share real-time states and updates, such as showing what they are doing in the moment, fostering a sense of immediacy and connection. Participants described the importance of sharing multisensory or emotional memories, imagining ways to let others experience meaningful moments, like a special memory or sensory detail, in richer and more embodied ways than existing media allows.
  4. (4)
    Granular, Intuitive, and Playful Privacy Mechanisms. Participants emphasized the need for privacy that feels natural and understandable, not overwhelming or rigid. They described space-based privacy, where the type of space, such as a public plaza or a personal bedroom, intuitively signals boundaries. They also described playful privacy mechanisms, such as secret knocks or hidden rooms, that make managing access feel engaging and enjoyable. Participants expressed the importance of contextual privacy, with the ability to adjust what different people see in different contexts or moments. They wanted to feel fully in control of their visibility, including the ability to become invisible or undiscoverable to people they block or want to avoid.
  5. (5)
    Refocusing Interaction Priorities for Meaningful Engagement. Participants expressed frustration with how mainstream platforms prioritize endless content streams, celebrity updates, and commercial content over personal connections with people they care about. They described the value of intentional content consumption, where they could choose what they wanted to engage with rather than being fed endless streams. They also advocated for reducing celebrity and commercial influence, preferring to focus on posts from close connections. Some participants described using spatial representations of relationships, such as placing their closest friends’ spaces physically nearer to them in the platform’s landscape. Across these ideas, participants emphasized the value of purposeful, meaningful use of time, where interaction feels worthwhile because it supports relationships, not just engagement metrics or passive scrolling.
These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams. Instead, youth express a clear desire for platforms that prioritize meaningful connection over superficial engagement. While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
Our participants envision environments where relationships deepen organically through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement, challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
5.2. Spatial Integrity for Meaningful Social Connections in Online Environments
Concept Description Supported Interactions
Spatial Presence The ability to perceive and navigate a virtual space, including awareness of other movement, and presence. Gaming and VR platforms excel at this, whereas mainstream social media provides limited spatial affordances for awareness and intentional navigation. Immersion and embodiment (Section 4.1); memory and experience sharing (Section 4.4.3); co-existing in surreal landscapes (Section 4.6.3)
Spatial Configuration The presence of fixed/defined spatial structures that facilitate interaction. Platforms like Discord create stable communal spaces with clear channels and streams. Users value this for fostering deeper relationships. Neighborhood structures (Section 4.3.1); prioritizing closer connections (Section 4.3.3); creating quality conversational spaces (Section 4.3.4); imposing clearer boundaries (Section 4.3.5); deliberate navigation and content discovery (Section 4.5.1)
Spatial Composition The segmentation of a social network into defined areas and pathways. Clear boundaries help manage social norms and interactions, with current platforms blurring personal and public spaces. Building comfort through ambient presence (Section 4.2.2); surfacing third places (Section 4.3.2)
Spatial Depth The amount of layered, rich information and space supporting richer, more intuitive posting and commenting. 3D social media limits expression to explicit actions like posting and commenting. Leveraging spatial cues for intuitive privacy control (Section 4.1.2); indirect self-expression through customization and artifact display (Section 4.4.1); memory and experience sharing (Section 4.4.3)
Table 2. Dimensions of Spatial Integrity in sociotechnical systems in supporting meaningful social connections.
Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces:
  1. (1)
    Spatial Presence: Understanding the Space and Who’s There. Spatial presence refers to the ability to perceive and navigate within a virtual space, encompassing elements such as proximity/propinquity (knowing who is nearby), navigation (moving within the space), and awareness (understanding one’s position relative to others). While gaming platforms and social VR environments often excel at creating this sense of presence, mainstream platforms like Instagram lack basic spatial affordances, limiting users’ awareness and intentional navigation within the platform.
  2. (2)
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offer only transient content streams. Our participants emphasized the importance of dedicated areas like communal lounges or study spaces where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
  3. (3)
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity; a user’s “feed” is meant to be a personal space yet often lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
  4. (4)
    Spatial Depth: How Much Information Can Be Perceived. Spatial depth determines how much information users can perceive in a space. 3D spaces enable higher perceptual density, allowing users to process information more intuitively (Kumar and Benbasat, 2004; McIntire et al., 2014; Tavanti and Lind, 2001). Current platforms, limited by 2D interfaces, restrict users to explicit actions like posting and commenting rather than supporting more nuanced forms of expression. Incorporating spatial depth could enable richer self-expression and more natural interactions.
According to our participants, these dimensions serve as the foundation for designing online spaces that effectively support the kind of social connections they desire. The relationship between these dimensions and the interactions they support is laid out in Table 2.
These four dimensions of spatial integrity—presence, composition, configuration, and depth—provide a framework for designing online spaces that support meaningful social connections. By prioritizing these dimensions, platforms can create environments where the kind of meaningful interactions our participants desire can flourish. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram. Especially as spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
5.3. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
There are societal concerns about social media’s impact on youth well-being (Oth, 2023). Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality. Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023). At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024). This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connection that youth actually seek. Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
Our interviews reveal how mainstream social media’s design fixations (Jansson and Smith, 1991) have shaped youth’s perceptions of what social platforms can be. Youth have come to equate social media with certain limiting characteristics: endless content streams, superficial interactions, and 2D interfaces focused on broadcasting posts. Interestingly, while they report meaningful social connections and positive experiences in other online social spaces like Minecraft or Discord, they view mainstream social media as inherently tied to emotionally disconnected experiences and superficial interactions. This cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The Hogwarts-inspired context provided fresh metaphors for exploring social spaces, privacy, and connection. Unlike traditional design workshops where youth typically propose incremental improvements to existing platforms, these sessions enabled participants to imagine entirely new possibilities for meaningful connection. More importantly, the workshops helped participants expand their understanding of what social platforms could be, empowering them to articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This study demonstrates how design methods can offer a generative path toward hope in domains where narratives of doom are prevalent.
5.4. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently generalize to youth who are not familiar with such environments. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research should empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him, including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.
6. Conclusion
This study explored how 3D gamer youth reimagine social media when given the freedom to think beyond the limits of today’s platforms. Through the fictional world of Hogwarts, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption that defines mainstream social media.
Their visions called for platforms that feel more like 3D games than archetypal ones like Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to the complexity of real relationships and incorporate elements of playfulness and whimsy.
Building on these insights, we introduced the concept of spatial integrity, a set of foundational design principles that make such visions possible. While spatial integrity alone does not guarantee meaningful connection, it provides the structural conditions that support the kinds of relational, embodied, and playful interactions participants described.
The study underscores that youth are not necessarily seeking less time on social media, but rather more fulfilling and emotionally present interactions within these digital spaces. Their visions point towards a future where social media feels more like a dynamic, shared social world like those found in 3D gaming environments. While our findings are rooted in the experiences of youth already familiar with 3D gaming, they offer a hopeful design direction for spaces that support meaningful social connections that youth desire.
Appendix A FI Workshop Protocol
Introduction
Welcome, and thank you for participating in our design workshop. Before we begin, I want to ensure that you’ve read the consent form I sent earlier. Have you had a chance to review it?
[Response]
Great! Just to reiterate, this session will be recorded for research purposes. Do I have your consent to proceed with the recording?
[Response]
Thank you. If at any point you feel uncomfortable or wish to stop the interview, please let me know.
General
Let’s start with some general questions about your experiences with 3D gaming and Harry Potter.
  • Which 3D games do you play and how long have you been playing them?
  • (Which is your favorite?) Can you describe the 3D elements or the spatial elements in that game a little bit, please?
  • Have you ever played in VR mode if that is available?
  • Any AR games you have played? Like Pokémon Go?
  • What Harry Potter books or movies have you watched?
  • Which is your favorite?
  • Which Hogwarts house would you put yourself in?
Narrative/Plot
As mentioned in the consent form, our goal today is to explore innovative ways to improve social media. We particularly invited 3D game players with a love for Harry Potter because we believe this will help us think creatively and “outside the box” about what an ideal, magical social media could look like, beyond the confines of a small 2D smartphone screen.
Don’t worry about being too creative—we’re here to guide you through the design process, not to evaluate you. We want to explore different design directions with your help, and we’ll be actively involved while also giving you the space to share your ideas.
To give our brainstorming some structure, we’ll start with a specific scenario to solve during this session. I’m going to ask you some guiding questions to help you address the scenario:
Imagine you are a student at Hogwarts. You have friends who are Muggle-born, friends from magical families, and even friends at other wizarding schools around the world. You’re also getting to know me, as I’m a new student you’ve just met at Hogwarts. Keeping in touch with everyone is tough because traditional Muggle social media like Instagram doesn’t really portray your wizard self very well. You don’t have a smartphone in the first place because Hogwarts people don’t need electronic devices, so you can’t really charge your smartphone.
One day, a brilliant Muggle-born student had an idea. They decided to use magic to create a magical social media where everyone—Muggles, Hogwarts students, professors, and friends from other magical schools—could all communicate and share their lives in the most ideal way. Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?
Character Map
Before we dive into the designing part, let’s first map out your relationships in your new life as a Hogwarts student. Use the Miro board link (Figure 1) to write down the names or nicknames of people that you would like to put in each section. For example, you might want to put “younger sister” or “Jane” under Family, some of your coworkers under Quidditch Team members, yourself under the house that you think you belong to, your older middle school friends as Muggles, etc.
Please note the two boundaries in the four rectangles for each house. You can place your closer friends in the inner box and not-so-close ones in the outer box. And you can put the same name in multiple boxes (or parallelograms) as you wish.
  • How do you currently communicate with or stay connected with each of these people via Muggle social media?
  • What are your biggest pet peeves/issues with Muggle social media that you’d like to solve with magical powers?
  • What moments do you feel bad/guilty about using social media, if at all?
  • When does time spent on social media feel meaningless/meaningful/fulfilling, if at all?
  • Pet peeves around privacy settings?
  • When do you feel most connected to your friends when interacting with them on social media?
  • What would a better social media be if you were to use adjectives to describe it? What does “better” mean to you?
Design
Now, let’s think about what this magical world social media might look like. Remember, there are no right or wrong answers here. We’re looking for your imagination and creativity, so feel free to think outside the box and have fun with it!
  • What key capabilities would this thing have for communicating with [GROUP]*?
  • If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)
  • What would you want to share about yourself on this ideal magical “thing” with [GROUP]?
  • If you had [GROUP’] join this “thing,” how might its form adapt to include this new group?
*: Iterate for 1) Romantic Relationships and Closest Friends, 2) Closer Hogwarts Friends, 3) Muggle Friends, 4) Friends at Other Magical Schools, 5) Quidditch Team Members, 6) House Elf
Connecting Back
  • What do you like about this magical “thing” compared to Muggle social media?
  • What do you dislike about it?
  • Imagine using the magical “thing” we discussed earlier. How do you think it could resolve some of the issues people might face with Muggle social media?
  • How might this new magical “thing” better support your needs for self-presentation and identity management compared to traditional social media?
  • What challenges might arise in using such a “thing” at Hogwarts?
Thank you so much for your time and insights. Your contributions are incredibly valuable to our research. Do you have any final thoughts or questions before we conclude?
Appendix B Participant Interview Summary
Table 3. Key Features for Envisioned Social Media from Participant Interviews
PID Envisioned Social Media Features
P01 Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates, pinned messages, and casual interactions; Mailbox updates; Public areas for open interaction & Interest-based rooms for community engagement; Area-based access control; Music & mood reflection in rooms; News hub or enchanted map for real-time global updates.
P02 Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
P03 Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; Floating holographic interface; VR-like navigation; Selective visibility & privacy settings; Muggle guest system in selective areas with limited abilities; Shopping district & immersive commerce; Immersive media; Customizable avatars & dynamic appearance.
P04 Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods and activities; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
P05 Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
P06 Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers like knocking on a bookshelf or moving objects in a set sequence; House elf companion (AI assistant); Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
P07 Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
P08 Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
P09 Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Customizable bulletin board for updates; Moderation via magical guardians; Room of Requirement-inspired group chats.
P10 Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
P11 Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
P12 Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; 3D interactive messaging & calls; Interest-based connection zones; Engagement-based prioritization.
P13 Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
P14 Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Mutual unfollowing & privacy indicators; Authentic avatars.
P15 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
P16 Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
P17 Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
P18 Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
P19 Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
P20 Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
P21 Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking, teleportation, and environment transformation; Immersive social events; Space-based access control.
P22 Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
P23 Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; "Hall of Fame" memory posts.