### v3

# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

Abstract.
We investigate youth visions for ideal remote social interactions, drawing on co-design interviews with 23 participants (aged 15-24) experienced with 3D gaming environments. Using a Fictional Inquiry (FI) method set in the Harry Potter universe, this research reveals that young people desire social media that functions more like immersive, navigable shared social spaces. Across these interviews, participants identified six key priorities for meaningful social connection over social media: intuitive social navigation, shared collaborative experiences, communal environments fostering close relationships, flexible self-presentation, intentional engagement, and playful social mechanics. We introduce the spatial integrity framework, a set of four interrelated design principles: spatial presence, spatial composition, spatial configuration, and spatial depth. Together, these principles outline how online spaces can be designed to feel more like meaningful environments, spaces where relationships can grow through shared presence, movement, and intentional interaction. Participants also described the FI process itself as meaningful, not only for generating new ideas but for empowering them to imagine and shape the future of social media.
social media; fictional inquiry; youth; design
conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3ccs: Human-centered computing Collaborative and social computing
1. Introduction
Meaningful social connections are critical for young people’s development and well-being (Hawkley and Cacioppo, 2010), and social media has great potential to support these. At the individual level, richer and healthier interactions support youth well-being and identity development (Kim et al., 2024c). At the societal level, platforms act as critical infrastructures for sustaining friendships and community ties, making their design choices consequential for social capital (Phua et al., 2017a; Ellison et al., 2007). In fact, many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n. d.]).
However, social media platforms increasingly shift away from facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025). While designers initially created these platforms around sharing personal updates with friends, they now emphasize entertainment, news, and engaging larger, less intimate audiences. Meta’s data illustrates this shift: between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025). Mark Zuckerberg described this change as aligned with the company’s focus on “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
This shift raises concerns about meaningful online relationships, particularly for young people who have grown up with social media as central to their social lives. Most young people engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023). Despite high engagement, many report dissatisfaction with their social interactions, not because they want reduced usage, but because they expect platforms to better support their connections (Kim et al., 2024c). Current designs prioritize passive content consumption over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
Given social connections’ importance for young people and social media’s vital role in facilitating them, we must explore new design possibilities. Young people frequently express dissatisfaction with current experiences, revealing a significant gap between their social needs and existing platform interactions. Mainstream platforms employ standardized design patterns that constrain creative possibilities and foster pessimism among users about social media’s potential. This creates a compelling need for innovative designs that complement existing platforms by emphasizing meaningful interpersonal connections. Therefore, we ask the following research question: What types of social interactions might current social media designs be overlooking or inadequately supporting?
To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) with 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n. d.]) and Roblox (Roblox Corporation, [n. d.]). Participants collaboratively imagined a social media platform designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This imaginative context allowed participants to step away from existing platform constraints and envision more varied and meaningful social interactions. Since researchers recognize 3D gaming environments as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated participants could leverage gaming insights to propose novel social media interactions. Specifically, our research addresses: What types of interactions do young people familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
Our findings reveal that young people with 3D gaming experience envision ideal social media platforms as immersive, navigable environments that designers should create to feel like neighborhoods, community centers, and shared social spaces. Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming. These visions highlight young people’s desire for social media that feels more relational, meaningful, and personally rewarding.
We contribute empirical findings on young people’s perceptions of current social media platforms’ shortcomings and their desired experiences. We propose the concept of spatial integrity for social interactions in online spaces, encompassing four dimensions of spatial affordances that facilitate meaningful social connections: spatial presence (understanding the space and identifying who is present), spatial composition (the space’s meaning and significance), spatial configuration (the space’s organizational structure), and spatial depth (the extent of perceivable information within the space). We demonstrate that the FI method in social media design contexts prompts creative idea generation, enables participants to transcend current design patterns, and empowers them to envision social media’s positive potential, leaving them more hopeful about future possibilities.
2. Related Work
2.1. Meaningful Social Connection in Online Environments
2.1.1. Defining Meaningful Social Interaction
Litt et al. (Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself,” involving positive emotions, information exchange, or concrete benefits.
The study identified three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services). Non-meaningful interactions lacked any such impact.
Strong ties (friends and family), community ties, and shared activities were associated with meaningfulness, but acted as facilitators enabling impact rather than primary sources of meaning. Other attributes linked to meaningfulness included advance planning, memorialization (photos or videos), and synchronicity. The communication medium itself—online or offline—was less important than these other characteristics.
We define meaningful social interaction as any interaction that participants would actively seek in an ideal scenario, reflecting their subjective meaningfulness. We expected these themes to emerge in participants’ descriptions of their experiences.
2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
Though social media platforms are primary connection tools for young people (Boyd, 2014a), many users report feeling lonely or disconnected despite widespread use (Twenge et al., 2019; Primack et al., 2017), highlighting a gap between intended purpose and actual experience.
While social connection remains a primary motivation for social media use, mainstream platforms like Instagram (All, [n. d.]), YouTube (you, [n. d.]), and Discord (discord page, [n. d.]) are designed for diverse user motivations including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022). From a uses and gratifications perspective (Katz et al., 1973), users pursue varied goals from building social capital (Phua et al., 2017b) to media consumption (Menon, 2022). Users seeking social connection may encounter experiences prioritizing other engagement forms, leaving social needs unmet.
This gap underscores the need for a clearer understanding of meaningful social connections in online environments. While social media presents risks like social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b) alongside benefits like marginalized group support (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and identity development opportunities (Davis, 2012; Lee et al., 2022), social connection remains central for many users. For young people increasingly relying on digital platforms, we must move beyond treating connection as one competing goal toward a focused understanding of how platforms can meaningfully support social relationships.
2.2. Social Technology Design Explorations in HCI
2.2.1. Designing for Meaningful Social Connections
HCI research has extensively explored designs fostering meaningful social interactions online at both platform and feature levels. Platforms like BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) introduced unique affordances emphasizing spontaneity, ephemerality, and creative self-expression. This research identified ways to build relationships through features prioritizing authenticity and playful interaction. However, users report increasingly toxic interactions as platforms gain popularity, with design choices encouraging network expansion and compulsive behaviors like tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), detracting from meaningful interaction goals.
Feature-level efforts include designs encouraging effortful communication (Liu et al., 2021; Zhang et al., 2022) and genuine connections (Stepanova et al., 2022). However, social media experiences are holistic, involving a complex interplay of multiple design aspects and user behaviors. Individual features alone cannot fundamentally alter overarching social media narratives.
Beyond archetypal social media, recent work investigates spatial social technologies emphasizing immersive, embodied interactions. Social VR platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide virtual spaces and embodiment for co-presence and shared experiences, facilitating more personal and tangible interactions than 2D environments. Platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend gaming and social networking, enabling virtual environment navigation, avatar customization, and dynamic group interactions. However, these platforms are not generally recognized as conventional social media.
What youth envision as ideal social media at the landscape and feature levels remains unclear, especially considering emerging technologies like spatial computing platforms. Understanding these perspectives is crucial for reimagining social media’s future to meet users’ evolving needs.
2.2.2. Video Games as a Source of Meaningful Social Connection
Video games are increasingly recognized as valuable environments for meaningful social connection, particularly among youth. Large-scale surveys show games serve as major venues for friendship development and maintenance, especially for boys (Lenhart, 2015c). Teens play with known friends and make new ones through online gameplay (Lenhart, 2015a), with over half reporting meeting at least one friend online (Lenhart, 2015a, b). These interactions extend beyond gameplay, supporting ongoing relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
Research highlights gaming’s social and emotional benefits, including cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014). Certain designs promote prosocial outcomes by encouraging players to help others in-game and beyond (Gentile et al., 2009). Games offer deeper, eudaimonic experiences involving meaning-making, personal growth, and connection (Daneels et al., 2023). These findings suggest video games can foster meaningful social connections among youth and highlight the potential for game-like features in social media.
2.3. Recognizing and Overcoming Social Media Design Fixation
Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented design challenge. Though generating divergent ideas is critical for innovation, we often impose unnecessary restrictions. This self-limiting process confines design to localized problems rather than exploring broader, optimal solutions.
Social media design faces similar fixation (Sundaram, [n. d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020). Ad-based revenue models incentivize increased time on apps (Oth, [n. d.]). Platforms use compulsive strategies like friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)), creating distrustful user communities (Kim et al., 2025). Rather than proactively exploring new paradigms of designs and interactions, much recent work on social media design has disproportionately focused on reactively addressing existing harm (Kim et al., 2024b). The domination of mainstream platforms along with this reactive approach to design leaves users dissatisfied and reinforces perceptions that all social media experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
Analogical transfer offers one powerful method for overcoming design fixation (Novick, 1988; Buchanan, 1992). Therefore, in this paper, we attempt to break out of the design and cognitive fixation that assumes the future of social media is bleak by freeing participants from real-world constraints and beliefs and leveraging elements from fictional settings to serve as sources for analogical transfer.
3. Method
We conducted co-design interviews with 23 participants aged 15-24 to examine how we might envision ideal social media platforms free from real-world constraints and design fixation on archetypal platforms. All participants had 3D gaming experience and Harry Potter (Rowling, 2015; har, [n. d.]) familiarity, a popular young adult fantasy series where some possess magical powers. We used the FI method to scaffold participants’ design idea generation. All procedures were approved by the university’s Institutional Review Board (IRB).
3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
FI is an exploratory co-design method enabling participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). FI creates environments where people imagine beyond present limitations through carefully crafted scenarios incorporating immersive elements like physical artifacts or role-playing activities. Participants take generative roles envisioning futuristic or fictional settings. The method effectively overcomes design fixation by encouraging analogies from fictional scenarios and free imagination outside real-world constraints. Unlike other ideation techniques, FI emphasizes developing futuristic design materials through collaborative designer-user dialogue (Dindler and Iversen, 2007).
Dindler and Iversen (Dindler and Iversen, 2007) outline three key implementation steps: defining the inquiry’s purpose (staging design situations, exploring future ideas, or driving organizational change); developing an appropriate narrative that participants find comfortable yet distinct enough from current practices to encourage new thinking; and creating compelling plots introducing tension or conflict to motivate workshop activities. Unlike traditional role-playing, FI encourages participants to maintain their identities and expertise while operating in fictional contexts, exploring new possibilities through personal motivations and capabilities while circumventing societal constraints.
We view video games as rich analogy sources for social media design, offering unique affordances and interaction patterns inspiring new approaches to meaningful social connections. By examining how games facilitate social interactions, we can identify adaptable design patterns for social media contexts, breaking free from fixation to explore novel alternatives.
3.2. Material: The Narrative Framework
Figure 1. Miro board mapping real-life relationships to Hogwarts-themed groups, designed to immerse participants in the Hogwarts-inspired FI narrative.
[…]
This is a visual chart designed to organize relationships and connections in the Harry Potter universe. There are three columns of labeled sections. In the middle, there are four sections, each representing one of the Hogwarts houses: Gryffindor (red), Hufflepuff (yellow), Slytherin (green), and Ravenclaw (blue). Each house box has its crest. To the left, there are two smaller boxes. The top-left box is labeled “Best Friends / Romantic Relationship(s),” featuring small images of Harry Potter, Hermione Granger, and Ron Weasley, with a space for a name or nickname. Below this, there is a smaller box labeled “House Elf,” featuring a picture of Dobby holding a glowing object and an area for a name. To the right, there are two more boxes. The top-right box is labeld “Family,” containing images of Sirius Black, James and Lily Potter, and the Dursley family. Below this, another box is labeld “Profs & Other Adults,” displaying images of Severus Snape, Dumbledore, Professor McGonagall, and Hagrid. At the bottom, there are three large sections. The leftmost section is labeled “Muggles” with a blank space. The middle section, labeld “Students from Other magical Schools,” displays crests from Beauxbatons, Durmstrang, Ilvermorny, and Castelobruxo. The rightmost section, labeled “Quidditch Team Members,” features an image of a Quidditch team in red robes. Each section contains an empty space for names or details to be added.
We selected the Harry Potter world as our FI narrative framework. The series follows an orphaned boy who discovers he is a wizard and attends “Hogwarts School of Witchcraft and Wizardry,” forming deep friendships while uncovering his destiny and clashing with the antagonist.
Among numerous magical worlds, Harry Potter proved particularly suitable for FI interviews. Its extensive popularity and cultural presence provide shared understanding, facilitating recruitment and engagement. By 2023, 98% of US youth aged 18-27 identified as casual or avid fans, with half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). The book series sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and was translated into over 80 languages by 2020 (The Rowling Library, 2020).
Set in modern England with magical elements like enchanted objects and spells, this context allows incorporating or discarding aspects of contemporary life, including social technology designs and constraints. Young protagonists and coming-of-age narrative align with our focus on youth identity and relationship building. Character transitions parallel real-world moves between schools or locations, enabling exploration of how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging suits examining youth identity expression and self-presentation in virtual spaces.
3.3. Procedure
The first author conducted all interviews over Zoom, with the second author joining one session to validate the FI method. Interviews averaged 90 minutes. We began by asking about 3D gaming experiences and Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (mug, [n. d.])) connecting across Hogwarts and Muggle worlds. We shared a Miro (mir, [n. d.]) board (Figure 1) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles). This encouraged thinking about relationships within the Harry Potter context and enhanced fictional design space immersion. We then asked about social media experiences, focusing on meaningful versus meaningless aspects and ideal platform visions.
During the design phase, participants described their ideal magical social media capabilities and what life aspects they would share. We maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines. On the other hand, we encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?” and “If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)” We concluded by asking participants to compare their envisioned Hogwarts social media with Muggle social media. The semi-structured protocol is in Appendix A.
3.4. Participants and Recruitment
We used purposive and convenience sampling for recruitment. The first and second authors shared recruitment materials (including screener survey links) on social media, while the first author distributed campus flyers.
Participants needed to be 15-24 years old per UN (United Nations, [n. d.]) and WHO (who, [n. d.]) youth definitions and consent to video/audio recording during 90-minute interviews. Video recording followed FI guidelines, enabling gesture analysis if needed. We included teenagers and young adults to capture perspectives from distinct life stages sharing focus on identity exploration, self-presentation, and relationships.
Participants needed sufficient Harry Potter familiarity (agreeing with “I remember some details about the plot, characters, and magical aspects”) and prior 3D gaming experience (e.g., Minecraft, Roblox). We did not require extensive knowledge of Harry Potter because we felt that would be an unnecessary bias that is not particularly helpful in envisioning new social media design. While this made it somewhat difficult for a majority of participants to remember specific artifacts from the Harry Potter world, they knew enough to recognize the complete and extensive freeing extent of the magical powers in that setting. Though we imposed no social media experience constraints, all participants had some social media engagement.
Table 1. Demographics Summary. Abbreviations: Familiarity with Harry Potter - Expert (“I remember every detail about the plot, characters, and magical aspects (like how the castle moves, spells, magical creatures, etc.).”), Proficient (“I remember most details …”), Familiar (“I remember some details …”). Social Media Platforms – IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace). Reasons for Using Social Media – Connect (Connecting with friends and family), Network (Networking for professional purposes), News (Staying updated with news and current events), Share (Sharing content (photos, videos, etc.)), Community (Joining communities or groups with shared interests).
PID Familiarity with HP Social Media Platforms Reasons for Using Social Media Gender Age Ethnicity
P01 Proficient IG, DC, BR, RD Connection, Network, Entertainment, Share Woman 19 Asian or Asian-American
P02 Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI Connection, Network, Entertainment, Share Woman 20 Asian or Asian-American
P03 Proficient IG, DC, TT, X, RD, TH, NP Connection, News, Entertainment, Share, Community Woman 18 Black or African-American
P04 Expert IG, DC, SC, TT, BR, X, RD Connection, Entertainment Woman 16 Asian or Asian-American
P05 Proficient IG, DC, BR, RD, FB, TH Connection, News, Entertainment, Community Woman 18 White
P06 Familiar IG, DC, SC, TT, RD, FB Connection, News, Entertainment, Community Non-binary 18 White
P07 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, News, Entertainment, Share Woman 17 Hispanic or Latino, White
P08 Proficient IG, DC, SC, RD, FB Connection, Entertainment, Share, Community Woman 19 White
P09 Proficient IG, DC, TT Connection, Network, Community Woman 21 Asian or Asian-American, White
P10 Proficient IG, DC, SC, TT, BR, RD, TH Connection, News, Entertainment, Share, Community Non-binary 17 Hispanic or Latino, White
P11 Familiar IG, DC, TT, X, RD, FB, TH Connection, News, Entertainment, Share Woman, Non-binary 19 Native American or Other Pacific Islander, White
P12 Familiar IG, DC, SC, TT, X Connection, Network, News, Entertainment, Share Man 19 Asian or Asian-American
P13 Familiar IG, DC, SC, TT, BR, X, RD, FB, TH Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P14 Expert IG, DC, SC, TT, BR, X, FB Connection, News, Entertainment Woman 18 Hispanic or Latino
P15 Proficient IG, DC, SC, BR, X, RD Network, News, Share, Community Woman 20 Asian or Asian-American
P16 Familiar IG, DC, SC, BR, X, RD Connection Man 18 Asian or Asian-American
P17 Expert IG, DC, SC, TT, X, RD, TH Connection, News, Entertainment, Share Woman, Non-binary 17 Hispanic or Latino, White
P18 Familiar IG, DC, SC, TT, X, FB Connection, News, Entertainment Woman 23 Asian or Asian-American
P19 Expert IG, DC, SC, TT, X, RD, FB Connection, Network, News, Entertainment, Share, Community Self-described, GenderFluid/Trans Man 19 White
P20 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, Share, Community Woman 17 Black or African-American
P21 Expert IG, DC, SC, RD Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P22 Familiar IG, DC, SC, TT, X, RD Connection, News, Entertainment, Share, Community Woman 19 White
P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
Eighty-four individuals completed the screener survey. We contacted 48 who met criteria; 23 provided consent and completed interviews. Mean age was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) women, 3 (13.0%) men, 2 (8.7%) non-binary, and 3 (13.0%) gender fluid participants. While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel, 2019), we note that our participant demographics skew toward women. Full demographic information appears in Table 1.
(Lenhart, 2015b, c)
3.5. Data Analysis
We analyzed interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides systematic yet adaptable qualitative data interpretation. Its independence from pre-existing theoretical frameworks suited our exploratory focus on youth-envisioned social media design through FI, enabling effective capture and analysis of diverse participant perspectives.
Since not all participants agreed to have their videos on during interviews, we did not conduct video analysis of gestures that participants sometimes used when describing their ideas. Additionally, drawing was not part of our protocol because we have experienced in prior interviews for other studies that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination. All data therefore consisted of transcripts from verbal interviews.
Analysis began with the first two authors independently coding the same two transcripts line-by-line using descriptive codes closely adhering to data. Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atl, 2024). After four iterations, we reached code saturation. The first author then coded all transcripts using the final code set. All code versions and the final codebook are in supplementary materials.
3.6. Ethical Considerations
Given J.K. Rowling’s controversial transgender statements (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in recruitment materials and surveys, reiterating our position during interviews.
4. Results
Drawing from their experiences with 3D games and familiar media worlds, participants imagined social interactions extending beyond mainstream platforms like Instagram. Through the speculative frame of exploring a “new magical ‘thing’ for social communication” for Hogwarts students, they envisioned how digital environments might enable more embodied, relational, and emotionally meaningful connection. Their visions converged around six interrelated themes that collectively contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs. We refer to this synthesized concept as “Social Media at Hogwarts,” or SMH, throughout this paper.
4.1. Intuitive Social Navigation
Youth sought interactions that feel natural, embodied, and socially meaningful, reflecting how people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy management in intuitive, socially familiar ways.
4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Participants described how embodied, immersive qualities like realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space fostering presence, emotional connection, and meaningful interaction. In contrast, current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation, explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
For others, fulfillment came from embodied physical actions. P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing,” explaining “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
This sense of presence was felt even in seemingly passive actions. As P01 explained,
“I don’t know, there’s just a feeling about it that makes it different. I can’t really explain. Depending on 2D and 3D, but I feel like in 2D, you can’t really change anything about their homepage. You just observe it. The most you can do is add a like or something, and that would change their like count. But then they would know who you are, so that takes away the anonymity. But if you’re just looking at it, then you’re not really affecting anything. That’s just for your knowledge. But in a 3D setting, even if you don’t leave any physical footprints, in a way, they’re still there because you opened the door, you went in, you sat down. Maybe just that mark in a way that you’ve left on the place takes away from the anonymity. I don’t know if that makes sense but.” (P01)
For P01, 3D interaction supported more embodied engagement and introduced a felt presence that lingered, even without direct interaction.
4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed using spatial and visual cues modeled after real-world environments to make managing access and privacy feel more intuitive, socially natural, and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
“I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home? If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive and natural. P11 suggested spatial contexts could reduce privacy control efforts, stating “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.2. Shared Experiences
Participants emphasized that shared experiences, through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
4.2.1. Fostering Shared Memories Through Collaborative Activities
Participants described shared activities as meaningful foundations for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” they described Roblox as fostering lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
The value of shared experiences extended beyond completing tasks. P21 described how walking could spark memorable conversations: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Participants saw low-pressure, ambient social interaction, like co-existing in shared spaces or listening to music together, as natural ways to build connections and shared memories. P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual study room with Lo-Fi music. These spaces would mirror Discord servers’ 24/7 activity, offering comforting presence even without direct conversation. Even “mindless scrolling” became meaningful when shared, as P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music often facilitated these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs,” reflecting on their experience:
“At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music. And it was one of the most awesome nights because we all knew in that moment that collectively we all enjoyed what we were hearing and it was just good. I don’t know how else to describe it. It was just nice.” (P19)
They added that music could serve as a “conversation starter” while providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
4.3. Communal Ambience
SMH was envisioned as a platform for intentional, engaging interactions with close friends, making it naturally less prone to social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features fostering organic, intentional, personalized interactions and closer friendships. They described several design features–neighborhood structures, shared gathering spaces, and stronger boundary controls–that could cultivate this communal, low-pressure environment.
4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
Participants wanted SMH’s 3D world to mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster social interaction and relationship-building. P22 described in detail:
“So, in a virtual world, you might have your house, and then you could go to a community board or town square with postings for different groups … From there, you could travel to those group spaces. I don’t know if you’ve ever played or seen Club Penguin, but in that world, you had your igloo to decorate and could visit the town … The town had different buildings for different activities. I’m imagining something like that, where you have your house where people can visit and talk with you, and if you want to meet new people, you can leave your house and go out into the town.” (P22)
Youth envisioned SMH reflecting real-world relationship dynamics, using physical distances within the landscape to represent social tie strength. Some participants envisioned SMH resembling a college campus or neighborhood, where physical closeness naturally fosters interaction. One participant elaborated:
“And in this magical world, I imagine the people around you in your neighborhood would be from the same house as you, kind of like how people live together in dorms on a college campus. The nearby stores would be the ones your group usually goes to. So you’d share that sense of togetherness because you all go to the same stores or belong to the same house. I think that would help build a real sense of community.” (P03)
P01 proposed a tiered layout where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be farther away, requiring a “little hike” to reach. This approach of aligning physical landscapes with emotional closeness would streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
Participants imagined SMH neighborhoods including shared third places where people could casually gather, lowering social barriers and creating connection opportunities. P05 described these as “neutral spaces” not tied to specific activities or relationships, where groups could gather for various purposes. They imagined “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
Participants described shared neighborhood spaces like common rooms, courtyards, or clubhouses where residents could casually interact outside personal spaces. P01 described gathering in a “courtyard or game room” at the neighborhood center. Drawing on Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities. P09 imagined a “virtual dining room” allowing casual conversation while eating together. Unlike Hogwarts’ segmented house-based accommodation, this dining space was open and inclusive, enabling cross-social group connections.
Others described public third places supporting casual, serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you.”
4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends and trusted communities. P19 criticized how platforms like Instagram have shifted away from their initial focus:
“We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page. We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference for close connections: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH being “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Participants described how themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
“Anyone you’re talking to in general, just emphasizing conversation between individuals more, rather than just posting on a thread. Like how I talked last time about Reddit, you know, there’s a thread, and then you make a comment on it, but there are a hundred other comments, so no one’s going to respond to you. I want to focus more on having actual conversations with people, making real connections, and more individualized chatting. I find it’s way easier to get to know people a lot better.” (P22)
Participants described how conversational space design could make SMH interactions feel safer, more personal, and more inviting than existing platforms. Many valued walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to broad, undefined audiences. P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations. They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
Participants also imagined personalizing these conversational spaces to make them feel more emotionally connected. P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories,” making even casual conversations feel more intimate. P09 expanded on this idea, imagining a “Room of Requirement”-like system (roo, [n. d.]) where users could create custom rooms on demand to match the “purpose” of their conversation. Participants also valued interest-based spaces, similar to Discord (discord page, [n. d.]), where a “common point” (P13) or “good talking point” (P07) could help break the ice and give conversations direction.
4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
To ensure SMH feels safe and personal, participants stressed needing comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools. P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (Huang and Vitak, 2022)) to people they wanted to avoid, based on mutual follows. They stressed needing stronger, more thorough blocking controls, including blocking “alternate accounts” and fully removing blocked person traces, explaining “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (Bojkov, 2019).
Participants wanted stronger control over their visibility, imagining features letting them opt out of being seen or discovered entirely. P03 suggested an “invisibility cloak” for moments when users “don’t want anyone to see you.”
Preventing unsolicited stranger contact was another priority. P22 proposed a Discord-inspired model where only people you’ve interacted with can contact you, while P13 argued people “should only be able to add you if you meet in person and exchange a QR code,” reflecting that “it’s too easy to find people on social media.”
4.4. Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This curated image culture creates “a lot of seriousness” (P23), reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting their full range without curation or performance pressures.
4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Participants consistently described customizable avatars and personal spaces as meaningful self-expression tools, offering far more creativity than MSM’s limited settings. P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
Customizable “3D avatar[s]” (P11) and profiles were key ways of showing individuality. P22 reflected on early Myspace (mys, [n. d.]), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames were ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display. P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right? So I think that’s part of you.”
Beyond visual customization, participants imagined SMH supporting ambient, indirect expression, offering subtle cues about what someone is doing, feeling, or thinking without formal posts. P01 wanted their avatar to reflect real-time states like current outfit or music, so friends could intuitively notice “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 imagined friends picking up on interest changes by observing virtual space updates over time, providing “an insight into… what I’m thinking.”
Several participants highlighted the contrast between 2D platforms’ flat, compressed feel and 3D environments’ immersive qualities. P16 noted that VR creates presence feeling more like meeting “real people… compared to profile pictures,” because you can pick up on “mannerisms” and details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
Participants described how these choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests. As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
Drawing parallels to the physical world, P13 pointed out how personal spaces like bedrooms carry a sense of intimacy and authenticity that translates well to SMH. They explained:
“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate. It’s where you spend most of your time, and it shows your personality the most. It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes or 30 seconds the way you can on Instagram. If you wanted to add someone there, all you’d have to do is remove a couple of things. It’s easier than physically changing a space.” (P13)
Finally, participants appreciated these ambient and low-pressure forms of expression for their lower social stakes. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling like they are performing for attention. They described this kind of expression as “more justified” and less “embarrassing” than posting content just to be seen by others.
4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships while reducing pressure to maintain a single, polished identity. Rather than being limited to one polished identity, as on MSM, participants wanted ways to represent their full range of interests, identities, and relationships. P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are. P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
4.4.3. Sharing Everyday Activities and Lived Experiences to Deepen Understanding and Connection
Participants valued sharing casual, everyday moments and rich, embodied memories, filling gaps between MSM’s typical highlights to create a more natural, meaningful extension of daily life and personal connection. P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants imagined SMH supporting personal, embodied memory-sharing to deepen understanding and reduce misinterpretation. P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n. d.]) as a way to create stronger emotional connections. Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully. P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
Some participants imagined these shared memories including sensory or emotional details beyond visuals. P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication. Others compared it to live streaming personal experiences. P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record. P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Participants emphasized that flexible, contextual privacy is essential for authentic self-presentation, allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented. They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life. P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
Participants also reflected on the challenge of managing different audiences and social contexts on MSM and imagined SMH supporting more seamless transitions between identities and spaces. P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts. Building on this, participants imagined contextual identity management, where users could shift how they present themselves based on the audience or situation. P03 proposed a model where everyone maintains a “base” identity but customizes additional layers. P22 similarly described having different avatars or profiles for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
Participants also stressed that contextual privacy should extend beyond people-based access to include temporal and relational factors. P14 proposed limiting access to different stages of one’s past identity, saying, “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” Others, like P21, suggested gradual openness, where people could selectively reveal more of themselves over time. As they described, “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.”
Participants also highlighted the importance of individual control over these transitions and boundaries. P14 explained that privacy settings should not be one-size-fits-all, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy as not just a technical feature, but as a core requirement for supporting the kinds of personal, meaningful, and low-pressure interactions youth hoped SMH would make possible.
4.5. Intentional Engagement
Participants described growing frustration with current social media platforms prioritizing passive consumption over meaningful connection. In response, they envisioned SMH as a space promoting intentional, socially rewarding interactions. This section explores themes reflecting their desire to refocus social media on authentic connection, purposeful engagement, and meaningful time use.
4.5.1. Encouraging Deliberate Navigation and Content Discovery
Participants imagined SMH encouraging deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with. P22 described frustration with Instagram, explaining “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage intentional, purposeful engagement when browsing content and navigating the virtual world.
P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful. Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
Participants imagined features prompting users to actively choose what they wanted to engage with before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was key to this intentionality. P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling. P02 preferred slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” While convenient features existed, participants valued walking and exploring, turning simple actions into shared social moments.
P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them. They described SMH as a world where you might “walk through a town to get there,” explaining, “It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
P05 believed this effortful navigation would naturally discourage superficial engagement: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond content navigation, participants emphasized shifting the platform’s core purpose from endless consumption toward meaningful social connection. They expressed frustration with current platforms promoting mindless, passive consumption and hoped SMH would foster meaningful, intentional connections through active social engagement. They described frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled. P14 captured this sentiment, describing spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
MSM’s infinite scrolling design contributed to this feeling. P11 explained “there’s always more content on social media,” fueling fear of missing out and urges to keep scrolling. P14 critiqued TikTok for exposing users to “clickbaity” low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
In contrast, participants described active, social engagement as far more rewarding. P16 shared that “anytime not scrolling” but “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 noted meaningful use comes when “actually talking to someone,” especially staying connected with people “not in close proximity.” They reflected that while users might spend more time on SMH, it would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to watching “50-second TikTok videos of somebody making slime for like an hour.”
Participants contrasted passive consumption with active participation in interactive environments. P10 explained that while doomscrolling feels unproductive, activities like Minecraft offer “progress” that can be “saved and come back to.” P06 noted they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
Reflecting on SMH’s potential, P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference. They explained SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting meaningful engagement, rather than passive consumption, ultimately mattered.
4.6. Playfulness and Whimsy
Participants described SMH as a platform feeling youthful, playful, and deeply immersive, offering entertainment plus opportunities for creative self-expression, social bonding, and emotional well-being. By embracing whimsy and delight, SMH would make social media feel more personal, less pressured, and more meaningful.
4.6.1. Embracing Creative Joy and Playful Self-Expression
Participants imagined SMH transforming social media into a space for creative joy and playful self-expression, offering freedom to experiment with identity, explore personal interests, and connect through creativity. P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
For some, personal customization felt like meaningful identity exploration. P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression. P22 added that SMH offered something they hadn’t found anywhere else, explaining, “I can really be creative here… I can make something that I haven’t made anywhere else, you know. That’s where the fun aspect comes in, really, for me.”
This creativity extends beyond individuals to the wider environment and community. Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa! They built that, that’s so cool” (P06). SMH felt alive and engaging whether users were socializing or spending time alone: “it’s not just like, Oh, the group chats are only fun when there’s other people here. No, it’s fun when we’re by ourselves, too, we get to do our own things” (P06). The interactive environment itself, where “you’re not just sitting there chatting, you’re also walking around the space… you’re interacting with whatever else is in that area” (P06), offered more to do than conventional platforms. Whether through building, exploring, or engaging with shared content, SMH transformed social interaction into something playful, creative, and deeply engaging.
4.6.2. Designing Engaging and Playful Privacy Mechanisms
Participants imagined privacy not as tedious and confusing but as part of playful experience, envisioning interactive, game-like mechanisms turning boundary-setting into something fun, discoverable, and socially meaningful. P11 envisioned controlling space access through gestures like “secret knock” or hidden objects revealing “secret passages” for trusted friends. Others suggested knowledge-based gates like “secret handshakes” (P21) shared among friends. Drawing from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces when correctly summoned.
P05 compared these features to “little Easter eggs,” remarking “there’s that little kid part of everyone” that makes such features enjoyable, contrasting them with mainstream social media’s “wall of settings” that often feel confusing.
4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities. P01 questioned the need for phones at all, envisioning wands projecting “holographic screens, like Jarvis in Iron Man,” or activating interfaces through spells. Participants also imagined alternative magical devices: VR glasses, a “makeup mirror” portal, or a “watch” that immerses users into SMH.
These magical affordances were not only imagined as novel interfaces but also as ways to make extraordinary experiences more accessible. P09 highlighted that SMH could give people “experiencing stress” or lacking access to “really, really beautiful landscape and environments” the chance to walk around and “admir[e]” calming, scenic spaces. P16 imagined that users could take part in activities they might never otherwise experience, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
4.7. Expanding Horizons
Given the speculative FI co-design workshop, SMH embodies youth aspirations and values for social media platform interactions. While acknowledging potential challenges like over-immersion or distraction, they viewed these downsides as worthwhile trade-offs for its benefits. Participants described SMH as an inspiring alternative to existing social media, offering self-expression opportunities and deeply personal, intentional connection. They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
4.7.1. Breaking Free from Platform Monotony with Novel Interactions
Participants saw SMH as offering better user experiences, departing from existing platforms’ repetitive structures, and inspiring hope for social media’s future. P06 appreciated how SMH combines desirable aspects: “You get friends who you can talk to, keep up with, and stay connected to. And you get things you can customize. You kind of get the whole package there.”
Participants appreciated SMH’s break from existing platform monotony. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. P17 envisioned SMH as “more interactive, kind of like if Discord and Instagram were combined,” enjoying Discord’s interactions with Instagram’s profiles and posting ability. P22 highlighted limited customization on other platforms, explaining “I just have a profile picture.” On SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” making the experience feel “novel,” like a “breath of fresh air.”
For some, SMH’s novelty lay in prioritizing enjoyment over seriousness. P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun. For example, P14 initially preferred a 2D version but shifted to 3D, reasoning:
“Yeah, [I would prefer the 3D version] actually 100%. I feel like,… the only reason I said 2D is because I felt like the need to be like a mature adult. But who cares? I feel like something that’s more enjoyable would be [the SMH].” (P14)
P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms. P06 observed that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” compared to “traditional ones.” The platform also evokes nostalgia and happiness. P15 remarked “Having this app would actually make them feel happy… with nostalgia,” anticipating SMH would be “a super big hit” with teenagers due to its playful, engaging nature.
4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
Through the FI process, participants realized they weren’t limited by existing platform structures. Instead, they felt empowered to imagine social technologies centering joy, creativity, and authentic human connection–qualities they had once believed beyond digital spaces’ reach. P10 pointed out that overcoming the design fixation around social media is crucial:
“The way that we can make social media better is if we expand the realm of what the definition of social media is. Instead of it being where someone just shares stuff, it can also be places where people connect, and people do this and that.” (P10)
Several participants noted that the FI workshop played a pivotal role in helping them gain a broader vision of social technologies and envision new possibilities for social media. P15 reflected on how their perspective evolved throughout the process:
“When we started, I literally was giving all these suggestions, which was completely the same thing which existed… I was repeating their concepts. So I think throughout this entire interview, I evolved my idea… actually thinking about what is new… So I think this has been a super hit. I like the entire conversation.” (P15)
They appreciated that SMH was not merely a “clone of the same thing… without actually adding anything else” but rather a concept that was “completely different… not even the same.” Similarly, P14 shared how the workshop shifted their initial skepticism:
“I don’t know what else you could do, because it’s online… so I really don’t know what else there could be to make it more meaningful.” (P14)
After engaging with the workshop, they felt empowered to imagine broader possibilities:
“I’m not really a tech person. So I don’t really think about [how to make social media better] that much. But I feel like this fake scenario… [helped] remind myself, ‘Girl, you have full control, this is no restrictions.’ After I got into that I… have a way bigger vision of what social media can become. (P14)
5. Discussion
Figure 2. 3D Gamer Youth’s Vision of Ideal Social Media Interactions
[…]
A Venn diagram comparing two categories of online platforms. The left circle represents platforms focused on social networking and broadcasting, with icons for Instagram, Facebook, and TikTok. The right circle represents platforms where youth engage in meaningful connections, with icons for Zoom, Minecraft, Discord, and Roblox. In the overlapping center, labeled “Social Media at Hogwarts (SMH)”, the diagram highlights youth’s vision of an ideal platform that combines the social visibility of networking platforms with the depth of meaningful connection found in gaming and community spaces.
Our analysis shows that youth imagine future social media not as a rejection of current platforms but as a reconfiguration of their strengths. They want the visibility and reach of social networking and broadcasting platforms while also craving the intimacy and authenticity found in spaces like games, voice chat, or small-group apps. Figure 2 illustrates this blended vision: an aspirational middle ground where meaningful connection is not an afterthought but the central purpose. Building on this insight, the following subsections trace how participants moved from critiquing existing platforms toward articulating design goals, culminating in our proposal of “spatial integrity” as a framework for supporting deeper forms of online connection.
The relevance of these findings has become increasingly apparent with recent technological developments. With the introduction of technologies like Genie 3 and similar platforms, the fictional world medium of remote social connection that participants envisioned is barely fictional. The 3D, open-world, and fully customizable platforms for self-expression and organic social connection that emerged from our workshops now represent concrete possibilities for the near future, much nearer than we anticipated when we began this study. This technological trajectory underscores the timeliness of our findings and their potential to guide the development of relationship-supporting technologies that leverage these emerging capabilities (DeepMind, 2025).
5.1. From Reactionary Designs to Aspirational Design Goals
There are societal concerns about social media’s impact on youth well-being (Oth, 2023). Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality. Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023). At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024). This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connections that youth actually seek. Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
However, in our earlier work with this youth population, we noticed several recurring challenges that shaped how participants thought about social media. First, there was a strong design fixation on social media design around existing apps, both in terms of interface, such as small 2D mobile screens, endless scrolling, and uniform templates, and in terms of user experience, where visibility depends on maintaining large networks of friends and followers, often at the expense of authenticity and comfort. Second, participants showed a cognitive fixation when imagining the future of social media, struggling to see it as anything beyond its currently perceived role as a convenient but limited tool for remote social connection. Third, these fixations were compounded by generally negative experiences with social media, which led many in earlier interviews to propose reactionary fixes, like stronger privacy, the removal of infinite scrolls, or eliminating general media content such as Reels. Such cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The procedure was intended to serve three purposes: breaking participants free from these biases, learned helplessness, and pessimism by encouraging them to think beyond the “app” or “platform” on mobile devices; equipping them with prompts that licensed thinking beyond perceived technical and practical constraints; and creating conditions where they could explore proactive and aspirational visions of social interaction rather than simply reacting to existing frustrations. Our chosen method prioritized uncovering underlying desires and needs that young people associate with meaningful remote social interaction over producing concrete feature lists.
In line with the goals of co-design processes (Sanders and Stappers, 2008), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most. Particularly noteworthy was the fact that the workshops helped participants expand their understanding of what social platforms could be, empowering them to feel more hopeful and articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This demonstrates how the design method chosen in our study offered a generative path toward hope in domains where narratives of doom are prevalent.
Our analysis contributes insights about the types of remote social interactions that youth find most meaningful and fulfilling, examining how these desired interactions align with or diverge from current platform affordances, as we detail in Section 5.2. We also identify core aspects of design that can undermine relational quality from which we develop the “spatial integrity” framework introduced in Section 5.3. Taken together, we hope this work equips designers with proactive design goals for envisioning new futures of social technologies and provides an initial framework for evaluating design integrity in relationship-supporting aspects.
5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings build on prior research showing that while youth are often dissatisfied with archetypal social media platforms, their goal is not to reduce time spent connecting online. Instead, they seek spaces that support deeper, more intentional social interactions. Participants described social connections that feel genuine, fulfilling, and emotionally present and not performative. The changes in social media platform design that they sought were as follows:
  1. (1)
    Immersive Experiences Through Spatial Elements. Youth expressed a strong desire to move beyond “flat” social media toward spatially immersive experiences. They imagined platforms where relationships could unfold through presence in shared 3D environments, not just posts on a feed. They described the value of feeling physically co-present with others, such as walking together in a virtual world or talking in a space that feels alive and embodied. They also envisioned personal spaces and community-centered landscapes, such as personal homes or neighborhoods that mirror real-world social structures. Several imagined futuristic forms of presence, such as those of holographic displays or VR interfaces that make interacting feel natural and embodied, reducing the sense of separation created by screens.
  2. (2)
    Organic and Intentional Social Interactions. Participants described meaningful connections as unfolding through casual, low-pressure, and community-oriented interactions. They valued flexible and safe spaces for direct conversations, where people could gather in protected and themed rooms to get to know each other more personally. They also emphasized the importance of community gathering places and shared activity spaces, such as areas to play games, collaborate on projects, or simply hang out. Visions of these spaces encompassed both low-intensity ambient co-presence, such as sitting together quietly while listening to music, and public third places that enable serendipitous encounters with new people. Participants also described the value of effortful and intentional navigation, preferring to move through spaces purposefully rather than having content or people pushed to them by algorithms.
  3. (3)
    Expressive and Lower-Stakes Sharing. Participants sought ways to express their identities naturally and casually, without the pressure of curating formal posts. They valued personal area and avatar customization as ways to reflect their personality through how they decorate spaces or design their avatars. They also envisioned ways to share real-time states and updates, such as showing what they are doing in the moment, fostering a sense of immediacy and connection. Participants described the importance of sharing multisensory or emotional memories, imagining ways to let others experience meaningful moments, like a special memory or sensory detail, in richer and more embodied ways than existing media allows.
  4. (4)
    Granular, Intuitive, and Playful Privacy Mechanisms. Participants emphasized the need for privacy that feels natural and understandable, not overwhelming or rigid. They described space-based privacy, where the type of space, such as a public plaza or a personal bedroom, intuitively signals boundaries. They also described playful privacy mechanisms, such as secret knocks or hidden rooms, that make managing access feel engaging and enjoyable. Participants expressed the importance of contextual privacy, with the ability to adjust what different people see in different contexts or moments. They wanted to feel fully in control of their visibility, including the ability to become invisible or undiscoverable to people they block or want to avoid.
  5. (5)
    Refocusing Interaction Priorities for Meaningful Engagement. Participants expressed frustration with how mainstream platforms prioritize endless content streams, celebrity updates, and commercial content over personal connections with people they care about. They described the value of intentional content consumption, where they could choose what they wanted to engage with rather than being fed endless streams. They also advocated for reducing celebrity and commercial influence, preferring to focus on posts from close connections. Some participants described using spatial representations of relationships, such as placing their closest friends’ spaces physically nearer to them in the platform’s landscape. Across these ideas, participants emphasized the value of purposeful, meaningful use of time, where interaction feels worthwhile because it supports relationships, not just engagement metrics or passive scrolling.
These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams. Instead, youth express a clear desire for platforms that prioritize meaningful connection over superficial engagement. While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
Our participants envision environments where relationships deepen organically through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement, challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
5.3. Spatial Integrity for Meaningful Social Connections in Online Environments
Concept Description Supported Interactions
Spatial Presence The ability to perceive and navigate a virtual space, including awareness of other movement, and presence. Gaming and VR platforms excel at this, whereas mainstream social media provides limited spatial affordances for awareness and intentional navigation. Immersion and embodiment (Section 4.1); memory and experience sharing (Section 4.4.3); co-existing in surreal landscapes (Section 4.6.3)
Spatial Configuration The presence of fixed/defined spatial structures that facilitate interaction. Platforms like Discord create stable communal spaces with clear channels and streams. Users value this for fostering deeper relationships. Neighborhood structures (Section 4.3.1); prioritizing closer connections (Section 4.3.3); creating quality conversational spaces (Section 4.3.4); imposing clearer boundaries (Section 4.3.5); deliberate navigation and content discovery (Section 4.5.1)
Spatial Composition The segmentation of a social network into defined areas and pathways. Clear boundaries help manage social norms and interactions, with current platforms blurring personal and public spaces. Building comfort through ambient presence (Section 4.2.2); surfacing third places (Section 4.3.2)
Spatial Depth The amount of layered, rich information and space supporting richer, more intuitive posting and commenting. 3D social media limits expression to explicit actions like posting and commenting. Leveraging spatial cues for intuitive privacy control (Section 4.1.2); indirect self-expression through customization and artifact display (Section 4.4.1); memory and experience sharing (Section 4.4.3)
Table 2. Dimensions of Spatial Integrity in sociotechnical systems in supporting meaningful social connections.
Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces:
  1. (1)
    Spatial Presence: Understanding the Space and Who’s There. Spatial presence refers to the ability to perceive and navigate within a virtual space, encompassing elements such as proximity/propinquity (knowing who is nearby), navigation (moving within the space), and awareness (understanding one’s position relative to others). While gaming platforms and social VR environments often excel at creating this sense of presence, mainstream platforms like Instagram lack basic spatial affordances, limiting users’ awareness and intentional navigation within the platform.
  2. (2)
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offer only transient content streams. Our participants emphasized the importance of dedicated areas like communal lounges or study spaces where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
  3. (3)
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity; a user’s “feed” is meant to be a personal space yet often lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
  4. (4)
    Spatial Depth: How Much Information Can Be Perceived. Spatial depth determines how much information users can perceive in a space. 3D spaces enable higher perceptual density, allowing users to process information more intuitively (Kumar and Benbasat, 2004; McIntire et al., 2014; Tavanti and Lind, 2001). Current platforms, limited by 2D interfaces, restrict users to explicit actions like posting and commenting rather than supporting more nuanced forms of expression. Incorporating spatial depth could enable richer self-expression and more natural interactions.
According to our participants, these dimensions serve as the foundation for designing online spaces that effectively support the kind of social connections they desire. The relationship between these dimensions and the interactions they support is laid out in Table 2.
These four dimensions of spatial integrity—presence, composition, configuration, and depth—provide a framework for designing online spaces that support meaningful social connections. By prioritizing these dimensions, platforms can create environments where the kind of meaningful interactions our participants desire can flourish. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram. Especially as spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
5.4. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently transfer to youth who are not familiar with such environments. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research should empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him, including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.
6. Conclusion
This study explored how youth with 3D gaming experience reimagine social media when freed from current platform constraints. Through Hogwarts’ fictional world, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption defining mainstream social media.
Their visions called for platforms that feel more like 3D games than Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to real relationships’ complexity while incorporating playfulness and whimsy.
Building on these insights, we introduced spatial integrity, foundational design principles enabling such visions. While spatial integrity alone cannot guarantee meaningful connection, it provides structural conditions supporting the relational, embodied, and playful interactions participants described.
The study underscores that youth seek not less time on social media, but more fulfilling and emotionally present interactions within digital spaces. Their visions point toward a future where social media feels like the dynamic, shared social worlds found in 3D gaming environments. While our findings stem from youth familiar with 3D gaming, they offer a hopeful design direction for spaces supporting the meaningful social connections youth desire.
Appendix A FI Workshop Protocol
Introduction
Welcome, and thank you for participating in our design workshop. Before we begin, I want to ensure that you’ve read the consent form I sent earlier. Have you had a chance to review it?
[Response]
Great! Just to reiterate, this session will be recorded for research purposes. Do I have your consent to proceed with the recording?
[Response]
Thank you. If at any point you feel uncomfortable or wish to stop the interview, please let me know.
General
Let’s start with some general questions about your experiences with 3D gaming and Harry Potter.
  • Which 3D games do you play and how long have you been playing them?
  • (Which is your favorite?) Can you describe the 3D elements or the spatial elements in that game a little bit, please?
  • Have you ever played in VR mode if that is available?
  • Any AR games you have played? Like Pokémon Go?
  • What Harry Potter books or movies have you watched?
  • Which is your favorite?
  • Which Hogwarts house would you put yourself in?
Narrative/Plot
As mentioned in the consent form, our goal today is to explore innovative ways to improve social media. We particularly invited 3D game players with a love for Harry Potter because we believe this will help us think creatively and “outside the box” about what an ideal, magical social media could look like, beyond the confines of a small 2D smartphone screen.
Don’t worry about being too creative—we’re here to guide you through the design process, not to evaluate you. We want to explore different design directions with your help, and we’ll be actively involved while also giving you the space to share your ideas.
To give our brainstorming some structure, we’ll start with a specific scenario to solve during this session. I’m going to ask you some guiding questions to help you address the scenario:
Imagine you are a student at Hogwarts. You have friends who are Muggle-born, friends from magical families, and even friends at other wizarding schools around the world. You’re also getting to know me, as I’m a new student you’ve just met at Hogwarts. Keeping in touch with everyone is tough because traditional Muggle social media like Instagram doesn’t really portray your wizard self very well. You don’t have a smartphone in the first place because Hogwarts people don’t need electronic devices, so you can’t really charge your smartphone.
One day, a brilliant Muggle-born student had an idea. They decided to use magic to create a magical social media where everyone—Muggles, Hogwarts students, professors, and friends from other magical schools—could all communicate and share their lives in the most ideal way. Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?
Character Map
Before we dive into the designing part, let’s first map out your relationships in your new life as a Hogwarts student. Use the Miro board link (Figure 1) to write down the names or nicknames of people that you would like to put in each section. For example, you might want to put “younger sister” or “Jane” under Family, some of your coworkers under Quidditch Team members, yourself under the house that you think you belong to, your older middle school friends as Muggles, etc.
Please note the two boundaries in the four rectangles for each house. You can place your closer friends in the inner box and not-so-close ones in the outer box. And you can put the same name in multiple boxes (or parallelograms) as you wish.
  • How do you currently communicate with or stay connected with each of these people via Muggle social media?
  • What are your biggest pet peeves/issues with Muggle social media that you’d like to solve with magical powers?
  • What moments do you feel bad/guilty about using social media, if at all?
  • When does time spent on social media feel meaningless/meaningful/fulfilling, if at all?
  • Pet peeves around privacy settings?
  • When do you feel most connected to your friends when interacting with them on social media?
  • What would a better social media be if you were to use adjectives to describe it? What does “better” mean to you?
Design
Now, let’s think about what this magical world social media might look like. Remember, there are no right or wrong answers here. We’re looking for your imagination and creativity, so feel free to think outside the box and have fun with it!
  • What key capabilities would this thing have for communicating with [GROUP]*?
  • If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)
  • What would you want to share about yourself on this ideal magical “thing” with [GROUP]?
  • If you had [GROUP’] join this “thing,” how might its form adapt to include this new group?
*: Iterate for 1) Romantic Relationships and Closest Friends, 2) Closer Hogwarts Friends, 3) Muggle Friends, 4) Friends at Other Magical Schools, 5) Quidditch Team Members, 6) House Elf
Connecting Back
  • What do you like about this magical “thing” compared to Muggle social media?
  • What do you dislike about it?
  • Imagine using the magical “thing” we discussed earlier. How do you think it could resolve some of the issues people might face with Muggle social media?
  • How might this new magical “thing” better support your needs for self-presentation and identity management compared to traditional social media?
  • What challenges might arise in using such a “thing” at Hogwarts?
Thank you so much for your time and insights. Your contributions are incredibly valuable to our research. Do you have any final thoughts or questions before we conclude?
Appendix B Participant Interview Summary
Table 3. Key Features for Envisioned Social Media from Participant Interviews
PID Envisioned Social Media Features
P01 Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates, pinned messages, and casual interactions; Mailbox updates; Public areas for open interaction & Interest-based rooms for community engagement; Area-based access control; Music & mood reflection in rooms; News hub or enchanted map for real-time global updates.
P02 Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
P03 Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; Floating holographic interface; VR-like navigation; Selective visibility & privacy settings; Muggle guest system in selective areas with limited abilities; Shopping district & immersive commerce; Immersive media; Customizable avatars & dynamic appearance.
P04 Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods and activities; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
P05 Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
P06 Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers like knocking on a bookshelf or moving objects in a set sequence; House elf companion (AI assistant); Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
P07 Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
P08 Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
P09 Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Customizable bulletin board for updates; Moderation via magical guardians; Room of Requirement-inspired group chats.
P10 Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
P11 Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
P12 Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; 3D interactive messaging & calls; Interest-based connection zones; Engagement-based prioritization.
P13 Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
P14 Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Mutual unfollowing & privacy indicators; Authentic avatars.
P15 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
P16 Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
P17 Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
P18 Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
P19 Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
P20 Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
P21 Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking, teleportation, and environment transformation; Immersive social events; Space-based access control.
P22 Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
P23 Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; "Hall of Fame" memory posts.